### v2

# The NEMESIS Catalogue of Young Stellar Objects for the Orion Star Formation Complex

Abstract
Context. The past decade has seen a rise in the use of machine learning methods in the study of young stellar evolution. This trend has led to a growing need for a comprehensive database of young stellar objects (YSO) that goes beyond survey-specific biases and that can be employed for training, validation, and refining the physical interpretation of machine learning outcomes.
Aims. We aimed to review the literature focused on the Orion Star Formation complex (OSFC) to compile a thorough catalogue of previously identified YSO candidates in the region including the curation of observables relevant to probe their youth.
Methods. Starting from the NASA/ADS database, we assembled YSO candidates from more than 200 peer-reviewed publications targeting the OSFC. We collated data products relevant to the study of young stars into a dedicated catalogue, which was complemented with data from large photometric and spectroscopic surveys and in the Strasbourg astronomical Data Center. We also added significant value to the catalogue by homogeneously deriving YSO infrared classification labels and through a comprehensive curation of labels concerning sources’ multiplicity. Finally, we used a panchromatic approach to derive the probabilities that the candidate YSOs in our catalogue were contaminant extragalactic sources or giant stars.
Results. We present the NEMESIS catalogue of YSOs for the OSFC, which includes data collated for 27 879 sources covering the whole mass spectrum and the various stages of pre-Main Sequence evolution from protostars to discless young stars. The catalogue includes a large collection of panchromatic photometric data processed into spectral energy distributions, stellar parameters (, spectral types, , , RV), infrared classes, equivalent widths of emission lines related to YSOs accretion and star-disc interaction, and absorption lines such as lithium and lines related to source’s gravity, X-ray emission observables, photometric variability observables (e.g., variability periods and amplitudes), and multiplicity labels.
Key Words.:
Catalogs – Astronomical databases – Stars: pre-main sequence – Stars: protostars – Stars: statistics – stars: binaries:general1 Introduction
As astronomy dives into the era of Big Data, we are now faced with the challenges of reliable applications of artificial intelligence and machine learning techniques. A major obstacle for unbiased machine learning applications in the domain of star formation stems from the data available to validate results, which are often susceptible to small numbers or limited to the specifications of a few larger-scale surveys. To address the scarcity of polyvalent databases of young stars, we revisited the published literature targeting the Orion star-formation complex (OSFC) under the framework of the NEMESIS to build a reference sample of Young Stellar Objects (YSOs) for use in upcoming research. We hereby report the work for data curation and a general description of the NEMESIS catalogue of young stellar objects for the Orion Star Formation Complex.
NEMESIS (New Evolutionary Model for Early stages of Stars with Intelligent Systems) is an H2020 project111https://nemesis.konkoly.hu/ aiming at revisiting the star formation paradigm with the aid of Big Data and machine learning techniques. NEMESIS efforts included a new version of the Herschel/PACS point source catalogue (Marton et al., 2024a) with an associated Deep Neural Network approach for source removal (Madarász et al., 2025); the study of the 2D morphology of YSOs in photometric images using self-organising maps (Hernandez et al., 2024 submitted), the variability characterisation of an all-sky sample YSOs from Gaia DR3 (Mas et al., 2025), a compendium of the application of deep learning techniques for the identification of YSOs all sky (Marton et al., 2024b, 2025 in prep) with the three latter presenting applications of the present catalogue. Additionally, two follow-up studies added value to the present catalogue through the fitting of YSO’s spectral energy distributions (SEDs) to synthetic libraries based on radiative transfer models (Gezer et al., 2025), and the ranking of bona fide YSOs (Roquette et al. in prep. 2025).
The Orion Star Formation Complex (OSFC; Bally, 2008) was chosen as our target as it is the largest and most diverse nearby star formation region within 500 pc of the Sun, with tenths of thousands of YSOs distributed over almost 600 square degrees in the sky (see Fig. 1). The OSFC harbours populations that span all ages relevant to studying young stellar evolution ( Myr; Kounkel et al., 2018) and covers the whole stellar mass spectrum from high-mass stars to intermediate- and low-mass YSOs (e.g., Hillenbrand, 1997; Muench et al., 2008; O’Dell et al., 2008). Previous research reported relatively large samples of YSOs located in the OSFC, but these were susceptible to the methodological specificities and limitations inherent to certain instruments or focused on subregions. For example, studies carried out with the Spitzer space telescope (e.g., Megeath et al., 2012, 2016; Fang et al., 2013; Getman et al., 2018) collectively identified YSOs. However, these are limited by Spitzer’s reduced coverage of the OSFC and favour the detection of less evolved YSOs with significant contributions of a disc or envelope to their SEDs. Large-scale kinematic surveys based on Gaia (e.g. Zari et al., 2018) characterised a similar number of YSOs, but are conversely biased towards optically visible sources. Similarly, large-scale spectroscopic surveys (e.g., Da Rio et al., 2016; Kounkel et al., 2018; Briceño et al., 2019; Hernández et al., 2023) focussed on the brightest sources located in less crowded areas and are accordingly biased toward more evolved YSOs, which already emerged from their envelopes, being near-infrared bright and optically visible. With the NEMESIS Catalogue of YSOs for the OSFC, we propose to go beyond these limitations by compiling a cumulative source list from the ensemble of previous research and combining complementation from public datasets and modern data science techniques to and mitigate, as possible, the biases from previous surveys.
Our data curation (Sect. 2) started from a historical data compilation based on peer-reviewed articles focused on studying YSOs in the OSFC (Sect. 2.1). This historical compilation was then complemented with data from large photometric and spectroscopic surveys (Sect. 2.2 and 2.3) and from photometric data available at the Centre de données astronomiques de Strasbourg (CDS; Genova et al., 2000, Sect. 2.4). The motivations and criteria for the inclusion of sources in our compilation are fully discussed in Sect. 3, along with the justification of the data types collated into our catalogue. These include stellar parameters, SEDs, equivalent widths (EW) of emission lines and other spectroscopically derived features relevant to the study of YSOs, infrared (IR) classes related to YSOs’ evolutionary stage and information on their X-ray emission - the collated data are also summarised in App. D. We further added value to our catalogue by employing the curated SEDs to homogeneously derive IR classes for of the sources in the catalogue (Sect. 3.3.1) and by utilising a multifaceted approach to evaluate and multiplicity of sources (Sect. 4). Finally, in Sect. 5 we discuss the incidence of massive stars and contamination by other types of sources in our catalogue.
2 Data Curation
We reviewed the published literature targeting OSFC with the goal of building the largest catalogue of candidates and confirmed YSOs in the region. Our definition of the OSFC field encompasses 564 square degrees that cover the box between RA 74.2 and 92 deg and DEC and deg. We considered YSOs explicitly listed as members of the Monoceros R2 region (at RA, DEC: 91.94825, -6.37850 deg) to be outside our scope, although some of those were still indirectly included in our compilation. The sky distribution of 27,879 collected YSOs is shown in Fig. 1. Fig. 2 presents a schematic representation of the data curation workflow further described in this section.
2.1 Data mining strategy for historical data
The last comprehensive historical literature review for the OSFC dates back to the Handbook of Star Formation Regions (Reipurth, 2008a, b; Bally, 2008) - more than 15 years ago. However, that review did not provide a list of known YSOs in the region that could give us a head start. Hence, instead of starting from previous literature reviews, we focused on the entire body of peer-reviewed publications targeting the region, which have been indexed by the SAO/NASA Astrophysics Data System (Kurtz et al., 2000). This exercise was carried out with the objective of identifying and data-mining publications that focused on the identification and/or characterisation of YSOs in the OSFC. With our scope on building a census of sources where either a protostar or a pre-main sequence star is already observable, we are, therefore, roughly covering Class 0 to Class III YSOs. Although some cores and Herbig Haro (HH) objects may be collaterally included in our historical data compilation, those sources and others intimately related to the earlier stages of the star formation (like starless cores, jets, globules, etc.) are beyond our scope.
At the high stellar-mass end, our compilation included both higher-mass YSOs, such as Herbig Ae/Be, as well as O-type stars, which albeit no longer young within their own evolutionary path, are a youth indicator for a coeval population covering the full mass spectrum. Whenever possible, we have identified and flagged massive stars inferred to be beyond their zero-age main sequence (ZAMS) (see also Sect. 5.2.1).
2.1.1 Identification of Relevant Datasets through the NASA/ADS Database
The SAO/NASA Astrophysics Data System (NASA/ADS; Kurtz et al., 2000) is a digital library operated by the Smithsonian Astrophysical Observatory under a NASA grant, which concentrates bibliographic collections for astronomy and physics. We used a Python wrapper, the NASA/ADS Developer API (Casey et al., 2017), to query NASA/ADS for bibliographic references that targeted the OSFC. We started with a simple query for Orion to retrieve bibliographic entries. From there, we restricted our search to the Astronomy Database, excluded non-refereed entries prior to 2020, and refereed entries published before 2018, but never cited. After this initial filtering, we retrieved textual meta-data (titles, abstracts, and keywords) for bibliographic entries.
Next, we applied a “bag-of-words” method to process these bibliography entries. We started by merging titles, abstracts, and keywords for each bibliographic entry into a single text-based entry. We then employed Natural language processing (NLP) methods to tokenise, lemmatisate, merge multi-word terms into single tokens, remove stop words, and homogenise spelling. The processed text was then scored on the basis of the occurrence of contextual tokens provided in an input vocabulary. This input vocabulary included terms such as Young Stellar Object, Star Formation, Young, Equivalent Width, Emission Line, Line Profile, Pre Main Sequence, Accretion, H-alpha, T Tauri, Herbig, disc, protoplanetary, IR-excess and protostar along with alternative versions of these terms, including their abbreviations, synonyms, and derived words.
A subset of bibliographic entries was highly scored on the basis of this vocabulary. This subset was examined by a human who inspected abstracts and manuscripts, labelled entries by thematics, and ruled out bibliographic entries that were not relevant to our compilation. For example, we excluded purely theoretical studies, investigations of YSOs in other star-forming regions that employed previously published data in Orion as a comparison to their results, studies reusing previously published observational data without relevant added value quantities, publications based on purely qualitative data, studies focused on the interstellar medium or the Orion nebula itself, and studies focused purely on massive stars.
After this initial human-with-domain-knowledge filtering, we reduced our list to 1201 bibliographic entries with observational data deemed relevant to our data curation. Next, this bibliography was carefully read and the data and acquisition methods were summarised as possible. This process allowed us to further discard publications where: (i) the data actually provided were not relevant to the NEMESIS project; (ii) the tabulated data were not provided by authors; (iii) the data provided was redundant and introduced in another related publication, (iv) data products were limited to photometric catalogues for the OSFC field without explicit YSO identification, such as the data could be fully retrieved via cone-search using the VizieR-SED tool (see Section 2.4); or, finally, (v) the data were obtained several decades ago and where already included in previous data compilations in a suitable machine-readable format, hence also redundant. This process also included the identification of relevant meta-data within the text of the papers that were worth tabulating.
This phase of our project was finalised in January 2023, although some more recent publications were later identified and manually included in the compilation. The processing of the relevant bibliography also allowed us to identify a dozen other scientific papers that included relevant data for young stars in Orion but did not mention the higher-level keyword “Orion” in their title, abstract or keyword list, hence not being identified by our first query with NASA/ADS.
2.1.2 Data retrieval
From the 2000s onwards, a growing number of authors made data associated with their peer-reviewed publications at least partially available at CDS via the VizieR catalogue service (hereafter VizieR; Ochsenbein et al., 2000), making data collation very straightforward once the relevant references and tables were identified. However, the data relevant to this compilation were often absent or only partially uploaded to CDS. We therefore further employed one or more of the six strategies detailed below to extract data from scientific publications. These are briefly described in order of efficiency (and priority).
1.
Data were available via VizieR, or could be retrieved with widely used tools under table access protocol (TAP Dowler et al., 2019), for example using Astronomical Data Query Language (ADQL Osuna et al., 2008), TOPCAT (Tool for OPerations on Catalogues And Tables; Taylor, 2005) or astroquery (Ginsburg et al., 2019).
2.
The table was available in electronic format at the publisher’s website in some standard data format (csv, ascii, tbl, etc.).
3.
Tables were only available as displayed on the publisher’s website in HTML or LaTeX format. TableConvert (v 2.4.2) tools to convert them to a suitable machine-readable format.
4.
Tables were only available as published, as part of a text-based .pdf file. The Tabula Tool (Aristarán et al., v. 1.2.1) to convert it into a suitable format.
5.
Tables were only available as part of an image-based .pdf (common for older publications) or as an image format within the publication. In these cases, a .png or .jpeg image of the table was inputted to a free OCR tool (typically OnlineOCR). However, we noticed that OCR tools available to us at the time were prone to imprecision affecting decimal numbers and symbols such as ‘+’ and ‘-’. Hence, this method required a careful visual comparison between the input image table and the output ASCII table to correct for imprecisions.
6.
Relevant data were not available in tabular format and had to be interpreted and retrieved from the publication’s text.
Tab. LABEL:tab:papers_with_data summarises the methods for data retrieval employed in each of the 217 scientific publications with data collated into the NEMESIS YSO catalogue for the OSFC.
2.1.3 Data Integration
We processed the data included in the historical compilation to build a master catalogue in which each YSO-candidate previously identified in the literature is attributed a unique identifier. Sect. 3 provides more details of this processing. For each new literature reference integrated into our compilation, the new data set was matched to the master catalogue preferentially based on the identifier codes defined in previous studies. When unique identifiers were unavailable, we employed TOPCAT’s Pair-Match to join tables based on coordinate matching in the FK5 system. We used a 2” matching radius as the default. After testing different radius to match publications sharing common identifier codes, this was most often the best one. However, in cases where the source publication explicitly described larger PSF or astrometric precision, a larger radius of up to 10” was required.
At each step, unmatched sources were added as new sources and received appropriate indexing. Our unique identifier, namely NEMESIS_ID, follows integer numbers from 1 to 27915.
2.2 Data from large photometric surveys
In addition to the compilation of YSO candidates, we have also built a reference photometric database for the OSFC field based on a list of large-scale photometric surveys, including observations at 34 photometric bands and covering the range wavelength range 0.15–160 . This allowed us to build a regular baseline for our panchromatic compilation, helping us to understand our historical compilation’s detection limits and completeness. Furthermore, by retrieving data directly from these surveys’ archives and following their user documentation, we could guarantee that the catalogues were properly cleaned before they were integrated into our database. It also helped verify the reliability of the data retrieved using the VizieR-SED tool (Sect. 2.4). Unless stated otherwise, we used a 2” matching radius to identify counterparts in our catalogue.
- GALEX DR6+7:
The Galaxy Evolution Explorer (GALEX) provides UV data in two bands at FUV (0.15 ) and NUV (0.2 ). We retrieved data for DR6+7 (Bianchi et al., 2017) through VizieR, cleaned them to remove sources contaminated with artefacts.
- PanSTARRS1 DR2:
Data from the Panoramic Survey Telescope and Rapid Response System DR2 (Pan-STARRS; Chambers et al., 2016) was retrieved through the PS1 CasJobs SQL Server. PanSTARRS1 covers most of the sky above Declination in the bands gPS1 (0.48 ), rPS1 (0.62 ), iPS1 (0.75 ), zPS1(0.87 ), and yPS1 (0.96 ). We retrieved aperture photometric data from the stacked version of the survey and cleaned the dataset as recommended in Flewelling et al. (2020) to remove sources of poor quality, suspected duplicates, and keep only the best quality detections.
- Gaia DR3:
Gaia provides all-sky photometry in the bands GG (0.50 ), GBP (0.59 ), GRP (0.77 ). We retrieved both photometric and astrometric data for the OSFC field through the Gaia ESA Archive. We followed the recommendations in the Gaia DR3 release papers and employed the metrics defined by Riello et al. (2021) to correct for inconsistency between different passbands. Then, as recommended by Riello et al. (2021) and Fabricius et al. (2021), we limit the effects of brightness excess towards the fainter end of the passband by limiting our dataset to stars brighter than mag. Finally, we applied the saturation corrections proposed in appendix C.1 of Riello et al. (2021) for the brightest stars.
- 2MASS PSC:
Data from the Two Micron All Sky Survey (2MASS Skrutskie et al., 2006) Point Source Catalogue (PSC) is available all-sky and were retrieved via the NASA/IPAC Infrared Science Archive (IRSA). 2MASS/PSC provides photometry for the near-infrared bands J2M (1.24 ), H2M (1.66 ), and Ks,2M (2.16 ). We used the 2MASS photometric quality flag (|ph_qual) to clean up the catalogue. Only sources with quality A, B or C were kept in the main catalogue. Low SNR sources (ph_qual=D) and and upper limit detections (ph_qual=U) are kept, but as limit values.
- UKIDSS DR9:
Data from the United Kingdom Infrared Telescope (UKIRT) Infrared Deep Sky Survey (UKIDSS; Lawrence et al., 2007) DR9 were available for part of the OSFC as part of the Galactic Clusters Survey (GCS) and were retrieved through VizieR. The survey includes the infrared bands: ZU (0.88 ), YU (1.03 ), JU (1.25 ), HU (1.63 ) and KU (2.20 ). We have cleaned the catalogue for duplicates, sources flagged as probable noise.
- UHS DR2:
The UKIRT Hemisphere Survey (UHS; Dye et al., 2018) covered most of the northern part of the OSFC in the JU and KU bands. We have retrieved the data available from DR2 (Bruursema et al., 2023) through the WFCAM Science Archive ADQL service. We have cleaned the catalogue for duplicates, sources flagged as saturated or probable noise.
- VHS DR5:
The Vista Hemisphere Survey (McMahon et al., 2013) covered the southern part of the OSFC in the YV (1.02 m), JV (1.25 m), and Ks,V (2.14 m) bands. We retrieved data from DR5 through VizieR and cleaned for duplicates, sources flagged as probable noise.
- WISE:
The Wide-field Infrared Survey Explorer (WISE Wright et al., 2010) observed the whole sky in four mid-infrared bands: W1 (3.4 m), W2 (4.6 m), W3 (12m), and W4 (22m). For the W1 and W2 bands, we collected data from the CatWISE2020 Catalogue (Marocco et al., 2021), which provides photometry extracted from co-added images produced as part of the unWISE extension (Schlafly et al., 2019) based on all available observations from WISE and NEOWISE (post-cryogenic reactivation WISE mission Mainzer et al., 2014). For the W3 and W4 bands, no new data were acquired after the cryogenic mission, and the best data release to date is still the All-Sky Release Source Catalog (AllWISE Wright et al., 2019). We retrieved AllWISE data directly from the survey website222https://wise2.ipac.caltech.edu/docs/release/allwise/. As the WISE survey was not originally designed for the specific study of YSOs, the survey’s source extraction process was not optimised for the usual sky background in star-forming regions - an issue previously acknowledged and discussed by a number of authors (e.g. Koenig & Leisawitz, 2014; Marton et al., 2019). The direct use of the AllWISE catalogues for studying YSOs is thus prone to include numerous sources with spurious photometric measurements, especially in the two longer wavelengths. In App. A, we discuss how we built a Random Forest classifier trained on labelled image stamps of AllWISE W3 and W4 bands to identify and remove probable spurious detections from our catalogue.
- SEIP:
Data from the Spitzer Enhanced Imaging Products (SEIP SEIP, 2020) were retrieved via IRSA. The SEIP Source List catalogue was downloaded and processed according to the Spitzer Enhanced Imaging Products Explanatory Supplement (2013). This data release included photometry for the four IRAC (Fazio et al., 2004) bands I1 (3.6 ), I2 (4.5 ), I3 (5.8 ), and I4 (8.0 ), and for the MIPS (Rieke et al., 2004) band at M1 (24). We adopted the photometry for a 3.8” diameter aperture. We noticed that the complete cleaning recommendations in the Explanatory Supplement were too restrictive, especially for regions of the OSFC with an augmented sky background typical of young star-forming regions. Therefore, we limited our cleaning steps to removing extended and saturated sources.
- Herschel-PACS PSC 2.0:
We have also collected data from the Herschel-PACS Point Source Catalogue 2.0 (Marton et al., 2024a), which recently employed a hybrid strategy that combined classical source detection and machine learning techniques to provide an enhanced version of the Herschel/PACS Point Source Catalogue with much higher completeness levels than previous versions. This catalogue provides data in the three PACS bands P1 (70 ), P2 (100), and P3 (160). For this specific survey, we employed a 3” matching radius.
2.3 Data from large spectroscopic surveys
We have further complemented our catalogue with stellar parameters derived as part of a series of large-scale spectroscopic surveys briefly summarised here.
Gaia DR3:
Gaia DR3 includes spectroscopic data products based on low-resolution spectra from the Blue and Red Photometers (BP/RP spectra ; Andrae et al., 2023; De Angeli et al., 2023) and medium resolution spectra from the Radial Velocity Spectrometer (RVS; Recio-Blanco et al., 2023). Along with a series of stellar parameters, the former includes the emission line (Sect. 3.6) with pseudo-equivalent widths estimated as part of the Extended Stellar Parametrizer for Emission-Line Stars (ESP-ELS) (Creevey et al., 2023), and the latter includes the Ca II infrared triplet (Lanzafame et al., 2023).
- RAVE DR6:
The Radial Velocity Experiment (RAVE Steinmetz et al., 2020) observed medium-resolution spectra (R 7500) covering the Ca-triplet region. In its Data Release 6 it provides stellar parameters (, , metallicity and radial velocity) for 21 sources in our catalogue.
- LAMOST DR10:
The General Survey of the Large Sky Area Multi-Object Spectroscopic Telescope (LAMOST; Zhao et al., 2012) has been spectroscopically surveying half of the sky with low- (LRS: ) and medium- (MRS: ) resolution spectra. In its current public release, DR10, LAMOST data products relevant to our compilation in 4 of its released catalogues (
LRS_stellar,LRS_astellar,LRS_mstellar, andMRS_stellar). We have downloaded the v2.0 of these DR10 catalogues directly from LAMOST DR10 official webpage . This included LAMOST-derived parameters for 5,180 sources in our catalogue. - GALAH DR4:
The recent Galactic Archaeology with HERMES Survey Data Release 4 (GALAH DR4 Buder et al., 2024) included parameters (, radial velocity, , , , EWs for , , and Li) from high-resolution spectroscopy () for 2058 sources in our catalogue.
- Gaia-ESO Survey DR5:
- APOGEE DR17:
Although results from a dedicated sub-survey of APOGEE in Orion are included in our historical compilation (Cottle et al., 2018), we have further complemented our catalogue with stellar parameters derived with their ASPCAP pipeline in the context of APOGEE DR17 (Abdurro’uf et al., 2022), including stellar parameters for 9,429 sources in our compilation.
- ASCC-2.5 V3:
The All-sky Compiled Catalogue of 2.5 million stars, 3rd version (Kharchenko, 2001) which provides a large compilation of spectral types for bright stars and includes spectral types for 1,069 sources in our catalogue.
2.4 VizieR Photometry Viewer
We have further complemented our database with photometric data collected using the VizieR Photometry viewer. Powered by CDS, VizieR-SED is a tool intended for the visualisation of photometry around a given sky position. Hence, for an input coordinate or source name, it returns photometric observations “extracted around a sky position from photometry-enabled catalogues in VizieR” (VizieR Photometry viewer, ), where this photometry data can be exported in typical flux density units. The interpretation of the tool’s output as an SED is cautioned as CDS provides “no guarantee that all photometry points correspond to the target, especially for extended sources” (VizieR Photometry viewer, )333As part of our experience with VizieR-SED we second this recommendation and strongly advise against the indiscriminate use of SEDs retrieved with the tool for scientific purposes that require precise SEDs. One of the issues we identified can be traced back to the zero points adopted for magnitude-flux conversions in the tools’ back-end. The zero point values used can be recovered from the VizieR meta data tables METAfltr and METAphot at VizieR’s ReferenceDirectory. For example, 2MASS zero points adopted by CDS can be traced back to Pickles & Depagne (2010), and are offset in relation to official zero point values recommended by the survey’s documentation (Cohen et al., 2003), yielding a offset in the flux densities converted by VizieR.. Nevertheless, VizieR-SED provides a powerful tool for data mining the CDS database, and we have employed it to further search for relevant archival photometry for our sources. For that, we have followed their API’s recommendation to develop a Python script to query VizieR around the coordinates of each of our YSO candidates. We initially used a search radius of 5”. This wider search radius was chosen purposely, as it is prone to including contaminants from neighbouring sources, and we have taken advantage of this to flag sources likely susceptible to such contamination (see Sect. 4.3).
For the specific purpose of complementing the SEDs of sources in our catalogue (Sect. 3.2), we have reduced this radius to 2” to minimise contamination by neighbouring sources. We visually inspected a large number of SEDs produced with VizieR-SED data and compared them with SEDs built using photometry from our historical compilation (Sect. 2.1) and large photometric surveys (Sect. 2.2), which had their magnitude-flux conversions carefully carried out by ourselves using meta-data available in their original publication. This procedure allowed us to identify and exclude from the SEDs data from surveys showing large systematics compared to the larger body of data available. We have post-processed VizieR-SED outputs to remove data coming from all CDS tables already included in the compilation (i.e., from Sect. 2.1 and 2.2).
We also noticed large spreads in flux density in the SEDs generated by VizieR-SED, especially in the optical range 0.4-0.8 m. By further inspecting these data original tables, we verified that although a fraction of the spread remained unexplained, most of these could be traced back to either multi epoch surveys (YSOs are typically photometric variables as discussed in Sect. 3.8) or to surveys targeting extended sources and publishing photometry extracted for different aperture sizes (e.g. SDSS). Finally, we opted to avoid this wavelength range and kept only data in the ranges 0.1-0.4 m and 0.8-1000m, which seemed less affected by these issues.
We have also identified a significant number of duplicated data, introduced by scientific publications re-reporting photometric data from large surveys. For example, although we have removed all 2MASS official data release tables - as this data has already been included in our compilation in Sect. 2.2 - many sources had VizieR SEDs with dozens of 2MASS JHKs data points, most of which having the same flux values but originating from different CDS tables. Together with multi-epoch data, this type of duplicate is problematic for quantities derived from SED-fits, as in the added-value quantities discussed in Sect. 3.3.1, as it biases the fit results towards the filters with a larger number of photometric points. We have, thus, further processed the SEDs from VizieR-SED to aggregate photometric measurements reported under the same filter name by averaging their fluxes weighted by their uncertainties and propagating these uncertainties accordingly.
3 Criteria for inclusion in this compilation and description of historical data-types
Throughout our historical data curation (Sect. 2.1), we focused on sources in the OSFC field previously reported as young in the literature. We have adopted a positive evidence rule for inclusion in the catalogue, where a given source had to be reported as a YSO candidate by any but at least one of the types of study discussed in this section. Our approach thus prioritises completeness over purity. In this section, we summarise the main YSO identification approaches employed by the scientific publications included in our historical compilation. We note that a revision of the different criteria adopted in the literature was beyond our scope. However, we were still interested in collating data that could support the ranking sources as bona fide YSOs. Hence, in this section, we also discuss a series of data types collected for this purpose. The various data types included in the catalogue are also summarised in Apps. C and D.
3.1 Colour-magnitude and Hertzsprung-Russell diagrams
Newly formed stars are still contracting gravitationally and acquiring part of their final mass from their surroundings. Because they are still larger and colder than their Main Sequence (MS) counterparts, YSOs are located above ZAMS when placed in colour-magnitude (CMD) or Hertzsprung-Russell diagrams (HRD). Our compilation includes all sources reported in the literature as candidate YSOs from CMD selections. Barrado y Navascués et al. (2004) estimate that optical-near-IR (OIR) CMD YSO selections based solely on photometry suffer at least contamination by field stars. This contamination level can be greatly minimised when spectroscopic constraints for stellar parameters are available, enabling reliable transposition from observed CMDs into HRDs. Although we included all CMD-selected YSO candidates in this compilation, we also collated products from spectroscopic surveys that can help address the purity of our YSO candidate list. These data products are summarised in Tab. and discussed in the following sections.
3.1.1 Spectral Types
Our historical compilation yielded spectral types for 7,485 sources. With the complementation by large surveys in Sect. 2.3, a total of 11,497 sources have spectral types. We tabulated these spectral types in the MK system, with luminosity class, or peculiar spectral characteristics, provided as an extra flag when available. When multiple previous studies provided independent spectral type derivations, a list of values is reported.
3.1.2 Effective Temperatures
Our compilation included 73,808 measurements of effective temperature () for 17,589 sources collated from 42 publications. These measurements have been flagged according to the method used for derivation, where: SpT-PHOT indicates derivations based on standard tables of combined spectral-type and photometric data (e.g., Pecaut & Mamajek, 2013); SED indicates derivations based on SED-fitting to models (e.g. Bayo et al., 2008); SPEC indicates derivations based on fitting of observed spectra to spectra libraries (Kounkel et al., 2018); PHOT indicates derivations based on fit to photometric data using Markov Chain Monte Carlo or Bayesian methods (e.g. Da Rio et al., 2012). We did not include estimations based solely on CMD placement or colour/magnitude- conversions.
3.1.3 Bolometric values
CMD YSO selection is only possible for less embedded sources that have a significant portion of their radiation already visible. For more embedded sources, a more useful parameter space is the Bolometric luminosity-temperature (BLT) diagram (e.g., Chen et al., 1995). Hence, we also collected 700 derivations of bolometric temperatures () for 381 sources derived based on SED-fitting methods in 3 previous studies.
3.1.4 Attributes not included
The derivation of luminosities requires knowledge of the distance to the sources and the amount of interstellar extinction in its line of sight. Luminosities estimated prior to Gaia often carry large uncertainties as a result of the uncertainties behind distance estimations. Moreover, extinction estimations are often degenerate with the derivations. Thus, we opt to leave these two quantities out of our data collection because they are often biased by the type of data and methods available at the specific time of publication.
3.2 Spectral Energy distributions
As part of our historical compilation, we have collated photometric data covering the 0.1-1000 m range and processed it into SEDs. When these data were provided as magnitudes, their photometric systems were identified from their publications, and magnitudes were converted into flux densities. We have preferentially adopted filter profile information from the Spanish VO Filter Profile Service (SVO-FPS; Rodrigo et al., 2012; Rodrigo & Solano, 2020). When photometric system information could not be recovered from the publication or the telescope/instrument webpages or official publications, we adopted generic values in the Johnson-Cousin systems. A significant number of photometric tables out there include data re-reported from previous studies. We attempted to reduce duplicates by comparing magnitudes in the same filter collected for each source down to a precision of 0.01 and merging duplicates. In the present version of our database, we have not included limit values. For data related to the large photometric surveys described in Sect. 2.2 we have discarded data collected in the historical compilation and instead introduced our own processing of the photometry of these surveys. This greatly complemented the SEDs and guaranteed that of sources in our catalogue had at least 5 data points in their SEDs with an average of 22 datapoints per SED. Examples of the SEDs collected are shown in Fig. 3.
Next, we employed data obtained with the VizieR-SED tool to further complement the SEDs. As described in Sect. 2.4, these data have been post-processed to obtain averaged values. Data from this complement is also shown in Fig. 3, and this complementation helped increase the average number of data points per SED to 29, with 96% of SEDs having at least 5 data points. The typical number of points per SED and their incidence per wavelength range are illustrated in Fig. 4. A general description of the SED database is provided in Tab. . Additionally, flags with the counts of SED data-points per wavelength range are also provided and are aimed at facilitating the usage of the catalogue.
3.3 Infrared emission by circumstellar material
At the earliest phases of young stellar evolution, the dusty envelope surrounding YSOs obscures their radiation in the optical wavelengths, which is reemitted at longer infrared and radio wavelengths. Various YSO identification and characterisation methods stem from studying this reemitted radiation. Following the seminal work of Lada (1987), one of such methods is based on the derivation of the IR spectral index, , which gauges the amount of circumstellar material as the shape of the SEDs of YSOs in the IR (m):
Lada (1987) initially assigned YSOs into three classes associated with the evolution of the forming stars from core collapse to a pre-main sequence (PMS) phase: Class I sources were protostars still embedded in their natal material, Class II were disc-bearing PMS stars (often associated with accreting T Tauri stars), and Class III sources were PMS stars where the dusty circumstellar disc responsible for the IR-excess of YSOs has mostly dissipated. Later, Class 0 sources were included to account for the youngest deeply embedded YSOs (Andre et al., 1993), and the flat-spectrum class described sources with characteristics between Classes I and II (Greene et al., 1994). Examples of SEDs of those standard IR classes are shown in Fig. 3. With the advent of large-scale IR facilities (e.g., Spitzer: Werner et al. (2004), WISE (Wright et al., 2010) and Herschel (Pilbratt et al., 2010)), the increase and diversification of the available IR data have helped uncover physical aspects of the evolution of the YSO that could no longer be fully explained within Lada’s standard classification scheme, leading to the proposal of alternative classification schemes that include classes such as Transition discs (e.g. Hernández et al., 2007a; Luhman et al., 2008a; Fang et al., 2009; McClure et al., 2010; Dunham et al., 2014; Kim et al., 2016; Fang et al., 2013; Espaillat et al., 2014; Grant et al., 2018).
Rather than looking at SEDs, a range of IR classification schemes combines selection cuts within sets of CMDs and colour-colour diagrams into decision tree models to isolate the loci occupied by YSOs in these diagrams (Gutermuth et al., 2009; Kryukova et al., 2012; Megeath et al., 2012; Kryukova et al., 2012; Hernández et al., 2007a, 2009, 2010, 2014; Koenig et al., 2015; Cottle et al., 2018). These approaches cover the initial identification phase for a substantial portion of the YSO candidates featured in our data set. Although these techniques include cuts to exclude sources with colours and magnitudes consistent with active galactic nuclei (AGN), Star-Forming Galaxies, sources from the asymptotic giant branch (AGB), polycyclic aromatic hydrocarbon emission and knots of shock-heated emission, contamination by these sources is expected and further addressed in App. E.
All sources reported by studies applying infrared classification schemes to study YSOs in the OSFC field were included in our historical compilation. More details on the data types compiled are given in the next section and in Tab. . Although we have not compiled infrared excess measurements, we included a flag pointing to bibliographic references that include these attributes in the database.
3.3.1 Standard IR classification
Because a reported YSO class in any given IR classification was a criterion for including sources in this compilation, we have accordingly also collected such classes as part of our historical compilation. Nevertheless, the evergoing evolution of infrared classification schemes makes the collation and homogenisation of infrared classes derived by different authors an unpractical task. Instead, we employed the SED database described in Sect. 3.2 to homogeneously infer IR classes in the standard classification scheme. For that, we used the procedure described in (Hernandez et al., 2024 submitted) to fit Eq. (1) to SEDs and derive for each source using all data available within 2–24m. We were able to derive values for 25 799 sources (Fig. 5), with the requirement that at least two data points at least 2 m afar in wavelength existed within the 2–24m range. For assigning IR classes, we have followed the IR class definitions as employed by Großschedl et al. (2019), which is outlined in Tab. 1. Beyond this added value, we have compiled 16 716 IR classes for 5 379 sources from 30 publications. Fig. 6 shows a comparison between our own -index values and the 4 largest samples of -index values we found in our historical compilation.
| Class | range |
| 0/I | |
| flat spectrum | |
| II | |
| III thin disc | |
| III no disc |
Fig. 6 illustrates that data availability as a function of wavelength and methodological differences in the derivation of can affect the values derived, where the spreads observed can be traced back to distinct wavelength ranges employed: 3–8m from Koenig et al. (2015) 2–24m from Großschedl et al. (2019), 3–24m from Megeath et al. (2012), and 3–22m from Getman et al. (2017), whereas Großschedl et al. (2019) shows the closest equivalence to our values by covering a range similar to ours but with different specific data available. We stress that our indices were derived using all data in the wavelength range m as available. We required a minimum of two data points at least 2 m away in wavelength for derivation. As a result of this requirement, reported values were fitted using only data in the range m. As further discussed in App. B, this reduced data availability in the mid/far infrared may result in an excess of sources miss-identified as having a thin disc. To avoid such bias, we recommend using our in light of the provided data availability flags. In addition to the flags for SED sampling as a function of wavelength from Sect. 3.2, in Tab. B we also provide flags with the maximum and minimum wavelength available for derivation.
3.4 Lithium Depletion
Lithium is quickly depleted when parts of the stellar interior reach temperatures of MK (Soderblom, 2010). The location within the stellar interior where this takes place depends on the stellar internal structure and energy transport mechanism. In Solar-type stars, this will happen at the base of the convective zone. In the most massive fully convective stars, this takes place in their core, and in brown dwarfs below , the core will never get hot enough to destroy its Li (Soderblom, 2010). The timescale for Li depletion is estimated between 10-50 Myr for M-type stars, 20 and a few Myr for K-type stars, but it can be much longer than YSO evolution timescales for F and G-type stars (Jeffries, 2014). Nevertheless, theoretical investigations suggest that episodic accretion in young stars can produce young stellar objects with significantly augmented central temperatures compared to non-accreting counterparts at the same mass and age (Baraffe & Chabrier, 2010). This effect can severely enhance lithium depletion in accreting YSO. Therefore, lithium detection is an excellent spectroscopic proxy for confirming the youth of YSO candidates. All sources towards the field of the OSFC identified as young through their lithium abundance evidence were included in our catalogue. Lithium observables were available as the EWs of the Li I (6708Å) line or as Lithium Abundances, . 9,185 measurements were collated from 31 publications for 6,384 sources. More details on the data types collected are provided in Tab. 8.
3.5 Gravity
YSOs are still contracting onto the MS and have weaker surface gravity in comparison to MS stars. For comparison, while YSOs have , main sequence stars have , and K&M-type giants have (with MESA stellar evolution models as reference; Dotter, 2016). Gravity observables are thus a valuable asset in distinguishing YSOs against older sources with similar spectral type. Measurements of are typically obtained by fitting grids of synthetic spectra to observed spectra and are widely available (Da Rio et al., 2016; Kounkel et al., 2018; Yao et al., 2018; Jackson et al., 2020; Kos et al., 2021). In fact, more than 57k measurements were collated from 13 publications, with 12,313 sources having at least one such measurement. We note that Fig. 7 points to a degree of contamination by giants in our catalogue. This type of contamination is generally expected in surveys selecting YSOs purely from CMDs or colour-colour diagrams (as in Sect. 3.1 and Sect. 3.3). We further evaluate this contamination level in Sect. 5.2.2 and App. E.
Although widely available, derivations require medium to high-resolution spectra. A cheaper alternative is the measurement of equivalent widths of absorption lines known as surface gravity tracers (e.g., Reid et al., 1995; Slesnick et al., 2006; Schlieder et al., 2012). For example, the equivalent widths of the Na I doublet lines are expected to be 2 to 3 times stronger in cool M-dwarfs than in PMS stars younger than about 10 Myr (Schiavon et al., 1997). Among many atomic and molecular bands identified as surface gravity tracers, EW observations of the lines Na I, K I, TiO, and CaH 3 have previously been reported for YSOs in the OSFC. We collated 1,880 gravity-related EW measurements from 9 bibliographic references, with measurements available for 1,439 sources. Hence 13,065 sources (47% of our catalogue) had gravity-related quantities collected. The general distribution of the gravity observables is shown in Fig. 7 and data types are summarised in Tab. 9.
3.6 Spectroscopic signatures of accretion and outflows
Emission lines are a defining characteristic of accreting YSOs (Joy, 1945; Herbig, 1962). These lines are typically formed in outflows or infalling magnetic flows, being intimately related to the accretion process (Hamann, 1994; Hartigan et al., 1995; Hartmann et al., 1994; Muzerolle et al., 1998a). Extensive previous literature, including observational results and predictions from radiative transfer modelling, establishes the relationship between certain emission lines and the phenomenology behind the mass-accretion process in YSOs (e.g., Hamann & Persson, 1992; Muzerolle et al., 1998b, 2001; Bouvier et al., 2007; Lima et al., 2010; Kurosawa et al., 2011; Natta et al., 2014; Alcalá et al., 2014, 2017; Wilson et al., 2022).
From an observer perspective, in addition to helping identify young stars, measurements of equivalent widths of accretion-related emission lines allow quantifying the flux in excess due to accretion, enabling estimation of mass accretion rates (e.g. Natta et al., 2004; Sacco et al., 2008; Fang et al., 2009; Alcalá et al., 2017) and help identify bona fide YSOs. With a best-effort approach, we collated a large number of measurements of relevant EW of emission lines. Relevant spectral features are briefly discussed in this section if they were available for stars in the OSFC and included in this compilation. Although estimates of mass-accretion rates were also often available, these are heavily dependent on assumptions such as the source radius and distance, and were therefore excluded from this compilation. A summary of EW of emission lines collated can be found in Tab. and their distribution is illustrated in Fig. 8.
3.6.1 Hydrogen Lines
The strongest emission line observed in the spectra of optically visible YSOs is the H line (e.g. White & Basri, 2003). Although broad H is often associated with the accretion process, narrower H emission can also be attributed to the chromospheric activity of young stars (Martin et al., 1998). Hence, the association of H emission with the accretion process requires a threshold for distinction from chromospheric activity. Historically, many authors have applied thresholds with EW(H) at around 10 or 20 Å to identify Classical T Tauri stars (CTTS) and down to 5 Å to identify weak-line T Tauri stars (WTTS). As the diversity of the YSOs studied spectroscopically has grown over time, spectral type-dependent classification criteria were proposed (e.g. Martin et al., 1998). Currently, the most widely used criteria are the ones of White & Basri (2003) and Barrado y Navascués & Martín (2003). However, the increasing availability of large samples of EW(H) seems to indicate that spectral type dependent criteria for discerning CTTS from WTTS fail to identify low-accreting T Tauri stars at the end of their star-disc interaction phase (e.g. Briceño et al., 2019). This ever-going evolution of classification schemes also means that the labels for actively accreting CTTS and non-accreting WTTS provided in the literature depend on the classification scheme in use at the time of their publication. Therefore, we suppressed these labels from our compilation and focused on measurements of the H line. Alternatively to EW, some authors focused on the full width at 10% of the peak of the H emission profile, , which is proposed as an accretion diagnosis unbiased by the stellar spectral type (White & Basri, 2003). Although much less widely available, we also collated these quantities. Measurements for the H line, whether it is EW(H) or , are one of the most widely available spectroscopic data products, with 37,449 measurements collated from 49 publications for 21,244 sources - of the sample had at least one H measurement.
In the absence of optical spectroscopy around the H line, other hydrogen lines can be used to constrain accretion. We have collated 2,644 EWs from 10 bibliographic references for 2,501 sources, including observations for other Balmer series lines (H, H, and H11), lines from the Paschen series (Pa and Pa) and Brackett series (Br, Br11). Finally, the emission of molecular hydrogen, H2, is often interpreted as a signal for strong, bipolar jets or wind and can offer insight into Class 0 outflow activity (e.g. Laos et al., 2021).
3.6.2 Other lines
Beyond hydrogen lines, various emission lines in the spectra of YSOs can be traced back to inflow and outflow of material inherent to the star-disc interaction and accretion process in these objects. We have collated 2,849 EWs from 16 publications for 1,485 sources in our catalogue which are related to these processes. The general motivations for their inclusion in the compilation are discussed in the remainder of this section.
He I lines are a tracer of outflow phenomena and accretion in YSOs (Kozlova et al., 1995; Sacco et al., 2008). The He I line profile is sensitive to the kinematics of the stellar wind (Connelley & Greene, 2010). Hence, under certain conditions, He I can be related to chromospheric rather than accretion (e.g. Dupree et al., 1992). Nevertheless, He I has been suggested as a better tracer for identifying low-accreting stars that would otherwise be deemed non-accreting based on traditional classification schemes based solely on the emission of the H line (Thanathibodee et al., 2022).
Ca II near-IR triplet in emission is also characteristic of classical T Tauri stars (Hamann, 1994; Hillenbrand et al., 1998). In the youngest YSOs, Ca II is likely produced in the protostellar magnetospheric infall of gas (Muzerolle et al., 1998a) and it is considered a common emission feature for Class I YSOs (Azevedo et al., 2006; Connelley & Greene, 2010), having previously been used in association with Na I and CO to distinguish them from Class II and III sources (Greene & Lada, 1996). Although Ca lines in emission are often associated with accretion, these lines are instead observed in absorption for low-gravity sources (e.g., Reid et al., 1995).
3.6.3 Veiling
Optical and IR veiling is also a predominant characteristic of actively creating YSOs (Herbig, 1962; Basri & Bertout, 1989). Veiling is thought to arise from an excess flux originating from high-temperature material in the inner disc regions, the accretion funnel, and the accretion shock regions, which is responsible for diluting the photospheric spectral lines while enhancing the spectral continuum. Veiling at a given wavelength, , is typically quantified from a YSO spectrum as the creation between the excess flux to the stellar photospheric flux (e.g., Basri & Batalha, 1990; Folha & Emerson, 1999). 34,358 values were collected from 5 publications for 9,588 sources. These are also summarised in Tab. .
3.7 Kinematic methods
Due to the short timescales involved in the star formation process, observed groups of stars born from the same cloud are still moving together in relation to the galactic centre, presenting kinematic properties common to the parent cloud. Hence, groups of young coeval stars can be distinguished from foreground and background sources based on their common kinematic properties. Kinematic methods - based on variations in the star’s position in the sky - are thus widely used for selecting members of young coeval populations. There are two complementary techniques for studying stellar kinematics: the study of proper motions and the study of radial velocities. Today, reliable proper motion measurements are widely available due to Gaia, including many studies that evaluate the membership of Gaia sources in the OSFC. Members of the OSFC identified by Gaia kinematic studies are included in our catalogue if they were part of a study focused on the OSFC or in star-forming regions (e.g. Kim et al., 2019; Zari et al., 2018). Large-scale studies = typically based on machine learning, specialised in the identification of clustered populations all-sky (e.g. Hunt & Reffert, 2021) were considered beyond our scope. In addition, we also collated 60,294 radial velocity measurements for 13,721 sources from 33 publications. However, two caveats must be kept in mind. First, the inclusion of kinematic members of the OSFC will include its massive population (Sect. 5.2.1). Second, radial velocities are susceptible to variations due to, for example, binarity, and these should ideally be examined in the context of an associated Julian date, which is often not reported in the studies included in our historical compilation.
3.8 Variability
Since very early studies by Joy (1945), variability has been recognised as a defining characteristic of YSOs and has since been verified across the electromagnetic spectrum (e.g. Stelzer et al., 2007; Cody et al., 2014; Rebull et al., 2015; Venuti et al., 2015; Sousa et al., 2016; Roquette et al., 2020). The OSFC has been historically pivotal in understanding the variability of YSOs (e.g., Herbst et al., 1994; Carpenter et al., 2001). Accordingly, we have included YSO candidates identified in the literature based on their variability traced back to the physical phenomenology behind the evolution of YSOs.
3.8.1 Variability amplitudes
Variability amplitudes are the most widely reported variability proxy in the YSO literature. Although amplitudes have been tabulated by a number of variability studies, we note that these values are highly inhomogeneous and may be challenging to interpret as an ensemble. As an alternative, rather than collating literature values, we further added value to the catalogue by using Gaia DR3 data to added the variability degree of our sources.
Although Gaia DR3 light curves are only available for of sources in our catalogue, an alternative proxy for variability amplitude can be obtained from Gaia’s DR3 mean flux uncertainties through the (Mowlavi et al., 2021, see their eq. (2)), which could be calculated for 24,623 sources in our catalogue using parameters from the Gaia DR3 mean photometry.
3.8.2 Stellar rotation observables
One of the most common variability processes in YSOs is the rotational modulation spots in the stellar surface, be it cool spots induced by magnetic activity, or hot spots induced by accretion. Measurements of variability periods are thus often associated with YSOs spin rates (e.g. Serna et al., 2021).
Due to this association with stellar rotation, we have grouped variability periods collated with measurements under a table focussed on stellar rotation observables (see Tab. 13). 42,133 such observables were available, for 16,774 sources from 35 publications. However, we disclaim that all variability periods collected from the literature are reported in Tab. 13 under the variable name Per without discerning the possible physical origins for these periods. For example, eclipsing binaries and occultation by disc material outside the co-rotation radius can also explain periodic variability in YSOs.
3.9 X-ray emission
High levels of X-ray activity are observed in YSOs at all stages of evolution from Class I to the ZAMS (Feigelson & Montmerle, 1999; Preibisch et al., 2005). Additionally, since the ratio between the bolomeric and X-ray luminosities for YSOs is expected to be larger than field stars with (e.g., Feigelson & Montmerle, 1999; Preibisch & Feigelson, 2005; Getman et al., 2005a), X-ray observations have been established as a powerful tool for identifying YSOs. X-ray observations are especially powerful for identifying Class III, which typically lack most of the optical signatures of youth visible in Class I and II sources (e.g., Walter et al., 1988). We identified 19 publications that provided X-ray observations for 4,162 likely YSOs in the OSFC. Whenever available, we collected X-ray luminosities, observed and/or corrected fluxes, hardness ratio, and hydrogen column density . More details on the data collected are provided in Tab. .
4 Multiplicity labels and possible duplicates
We curated labels addressing the multiplicity of sources in our catalogue in four stages: through the historical compilation (1,539 sources labelled in Sect. 4.1), the use of all-sky binary catalogues from the literature (1,603 sources, Sec 4.2), the use of Gaia DR3 data (6,776 sources, Sect. 4.4) and derived using big data approaches (
6,912 sources, Sect. 4.3). As a result,
18,930 bin_type labels were assigned to
12,155 sources, which can assume the following values:
S - Binary identified from Spectroscopy: sources that were identified as spectroscopic binary (single or double-lined) or had observed variability in their radial velocity measurements associated with binarity.
E - Eclipsing Binary: sources that were reported as such in studies focused on their light curves.
A - Astrometric Binary: binaries identified based on their proper motion variations;
B - Generic: sources identified as a multiple system by previous studies being re-reported in a given study included in the compilation without enough accompanying information to be categorised elsewhere.
V - Visual pairs: source identified as multiple from surveys with high angular resolution444We note that the term “visual binary” is often used interchangeably with “visual pair”. The former term refers to physically associated double stars, which are observed close to each other in the sky as a result of their orbital movement. The latter refers to stars that appear close together in the sky, but are not necessarily physically associated..
U - Unresolved pairs: Unresolved close-pairs, typically indirectly identified using Gaia data (see Sect. 4.4)
Bl - blended sources: sources likely blended with a nearby source555Blended sources are sometimes observed as visual pairs and sometimes as unresolved pairs. (see Sect. 4.4).
R - RUWE unresolved candidate - Likely unresolved pairs identified using Gaia’s RUWE;(see Sect. 4.4.3)
? - candidate: sources reported as binary candidates because the detection criteria are close to the reported significance threshold.
Figure 9 presents the relative numbers of sources in each category. Some sources have been assigned to multiple categories. For example, 28% of the sources reported as eclipsing binaries in the literature were also reported as binaries from spectroscopic investigations. We note that we have not attempted any distinction between binary or triple systems.
4.1 Multiplicity labels in from the historical compilation
4.2 Multiplicity labels in from all-sky catalogues
We further searched the literature for large catalogues of binary stars that included the OSFC field and matched these to our source list. This helped us to attribute multiplicity labels to 1,735 sources, 1,102 of which were unlabelled in the historical compilation. This included:
(V?) 397 spatially resolved binaries with separations up to 200 AU identified with Gaia eDR3 data (El-Badry et al., 2021) with low chance alignment ;
(E) 11 eclipsing binaries identified using the first 2 years of TESS observations (Prša et al., 2022);
(S) 31 double-lined spectroscopic binaries identified by Kovalev et al. (2024) using values from spectral fits to LAMOST-MRS spectra;
(S?) 460 double-lined spectroscopic binaries identified by Zheng et al. (2023) using a deep-learning approach to analyse LAMOST-MRS spectra;
(S?) 36 spectroscopic binary candidates or RV-variables identified by Qian et al. (2019) due to their large radial velocity variations in their LAMOST-LRS spectra;
(S) 19 double-lined spectroscopic binaries identified by (Traven et al., 2020) with spectra from the GALAH survey;
(S?) 2 spectroscopic binary candidates identified by Birko et al. (2019) due to their radial velocity variability in the RAVE survey;
(S?) 7 spectroscopic binary identified by Jack (2019) from a sample combining Gaia DR2 RV and RAVE spectra;
(B?) 10 ellipsoidal variable candidates selected from their TESS light-curve variability (Green et al., 2023);
(B) 615 sources from the Identification List of Binaries Catalogue, which provided a survey of surveys of literature focusing on binary stars prior to 2015 (Malkov et al., 2016);
(S) 148 sources from GALAH DR4 (Buder et al., 2024) with a secondary component detected and for which primary and radial velocity could be measured.
4.3 Big data identification of visual-pair candidates
In Sect. 2.4, we have employed the VizieR-SED tool to retrieve the photometric data available on CDS around the position of each source in our catalogue. In this section, we further employ these data to gauge sources’ neighbourhoods.
Starting from the raw VizieR-SED output for a 5” searching radius, rather than relying again on the procedure described in Sect. 3.2 we carried a simpler preprocessing. We kept only data in the wavelength range 0.01-5, excluding multi-epoch surveys (e.g., TESS) and surveys with source detection and extraction at different apertures (e.g., SDSS). We then grouped the remaining data by unique CDS table name and photometric band and counted how many unique pairs of RA, Dec (down to a precision) existed. This process was repeated for each pair of CDS table name and photometric band, and the maximum number of unique sources inside the searching radius was stored as representative of the neighbourhood of that source. The same procedure was repeated using a 2” searching radius as well. The distribution of the number of neighbours for sources in our catalogue is shown in Fig. 10.
This procedure allowed us to leverage the large amount of photometric data available within CDS to investigate the neighbourhood of sources in our catalogue, including data observed by instrumental setups yielding a wide variety of pixel scales and PSF sizes, while being agnostic to the selection window introduced by the criteria for inclusion in this compilation discussed in Sect. 3. We found that 6,912 sources in our catalogue have at least one neighbour within 2” from the sources’ adopted coordinates, and 12,060 have neighbours within 5”. Sources with at least one neighbour within 2” were labelled as visual pair candidates (V?) with the support that 71% of these sources were also identified as such by other multiplicity labels considered.
4.4 Multiplicity labels in from Gaia DR3
The numerous avenues for investigating sources’ multiplicity using Gaia DR3 data merit this dedicated section.
4.4.1 Non-single source tables in Gaia DR3
We identified 223 sources with a counterpart in at least one of a series of tables focussed on non-single stars released as part of Gaia DR3 (Gaia Collaboration et al., 2023a), which are composed of sources identified as astrometric (150 sources from ; Halbwachs et al., 2023), spectroscopic (18 sources from ; Gosset et al., 2024) or eclipsing binaries (57 sources from ; Mowlavi et al., 2023), and with orbital models for all sources compatible with a two-body solution (10 sources from ; Holl et al., 2023b).
4.4.2 Gaia’s scan-angle effects:
Gaia’s readout window has a different size depending on the magnitude of the source and photometric band. For example, the typical window for a faint source in is , and for and (e.g., Riello et al., 2021; Holl et al., 2023a). In addition to this “rectangular” window, to achieve full-sky coverage, Gaia has a non-trivial scanning law, including a time-varying scan-angle. While for isolated point sources, these scan-angle variations are irrelevant, for close pairs, they are determinant factors between the source being resolved or blended. For example, pairs with separations will always be blended BP and RP while sources with separations will occasionally be resolved when observed close to along-scan angles and can therefore be identified using Gaia’s quality control flags. Meanwhile, pairs with separations are sometimes resolved with a six-parameter solution for the G-band (see also their Fig. 6 Lindegren et al., 2021).
Scan-angle-dependent signals:
Gaia’s time-varying scan-angle is hypothesised to induce scan-angle-dependent signals when observing unresolved binaries or extended sources. We followed (Holl et al., 2023a) procedure to identify 191 sources likely susceptible to scan-angle biases compatible with unresolved pair numerical models (separations mas), which we labelled as U? (candidate unresolved pairs).
Scan-angle-dependent spurious variability:
Gaia’s time-varying scan angle can also result in specific spurious periodic signals biasing Gaia’s epoch data. Holl et al. (2023a) and Distefano et al. (2023) attempted to identify such spurious periodicity by quantifying the correlations between Gaia’s epoch photometry and: (i) the image direction
parameter determination goodness of fit (), and (ii) the epoch-corrected excess factor (); with these parameters released as part of DR3’s table vari_spurious_signals. We identified 198 candidate unresolved sources (U?) by selecting sources with a correlation greater than 0.8 in both of these parameters.
Blended sources:
Due to Gaia’s several instruments and its different window sizes for the , and the and bands, sources resolved in the first may be blended in the others. Gaia DR3 data processing provides flags that attempt to quantify the number of transits in which a source was blended with a nearby source. Riello et al. (2021) employs these flags to define a metric, 666Following Riello et al. (2021), , where is the number of transits in the band (BP or RP) contributing to Gaia’s mean photometry in that band, which were flagged to be blends of more than one source inside the observing window; and is the total number of observations contributing that band’s mean photometry., to gauge the blending fraction of source. We used this metric to flag as Bl (blended) 3,028 sources that had more than 20% of their Gaia DR3 transits flagged as such ().
4.4.3 Gaia’s Renormalised Unit Weight Error
When observed by astrometric surveys like Gaia, stars with an unresolved or faint companion will show a different motion in their centre of light and centre of mass, which yields poor results when attempting to fit single-source astrometric models. In Gaia, such sources can be identified in terms of their Renormalised Unit Weight Error (RUWE), which is expected to be close to one for single sources with a well-defined centre of light and uniform motion (Lindegren, 2018). Since DR3, the threshold of (Lindegren, 2018) widely used with DR2 data to distinguish between single and multiple sources has been the subject of debate. For example, studies focused on eclipsing binaries have found a strong correlation between RUWE values and the separations of the unresolved binaries down to (see their fig. 3; Stassun & Torres, 2021). As part of the GaiaUnlimited777https://gaia-unlimited.org/, Castro-Ginard et al. (2024) provide a model to estimate suitable thresholds for binary selection with RUWE as a function of Gaia’s selection function variations as a function of sky position. This allowed us to choose appropriate RUWE values at the position of each source in our catalogue. We found an average threshold of ruwe for sources in our catalogue. GaiaUnlimited also provides tools to estimate the probability of erroneous selection of a single star with RUWE as a function of sky position, with a probability in the OSFC field. With 24,635 sources in our catalogue having at least one Gaia DR3 counterpart within 2”, and 23,716 sources having a ruwe measurement in DR3, 4,625 sources were flagged R based on our RUWE thresholds and should be interpreted as probable unresolved multiple systems.
4.4.4 Gaia’s RUWE in the context of a large sample of YSOs
As an illustrative application of our catalogue, we further contribute to probing previous suggestions that Gaia’s RUWE may be inflated in YSOs. As discussed in Sect. 4.4.3, Gaia’s astrometric solution assumes that sources are point-like with a well-defined centre of light and a single-star movement. Along with multiple systems, extended sources also deviate from these assumptions due to their non-point-like nature and the underlying larger uncertainty in resolving their centre of light. In YSOs, the diffuse emission of the envelope in Class 0/I sources, or even the radiation re-emitted by a thick disc in Class II sources could be the culprit of an uncertain centre of light.
Does the presence of a disc inflate Gaia’s RUWE?
The inflation of RUWE values due to the presence of a circumstellar disc has previously been suggested by Fitton et al. (2022), based on a modest sample of 122 single stars selected from adaptative optics surveys for which they find an excess of disc-bearing sources in the RUWE range 1–2.5. Here, we contribute to probing this evidence by examining a much larger sample of YSOs. For this end, we selected sources from our catalogue with both derived in Sect. 3.3.1, and Gaia’s DR3 RUWE.
We defined two subsamples based on (Sect. 3.3.1), namely sources with a thick disc () and discless sources (). To maximise the reliability of sample selections using the indices, we considered only sources with at least one datapoint available for m. We note that the requirement for Gaia data will inevitably remove the least evolved and most embedded Class I/O sources from our catalogue, as they are too faint and beyond Gaia’s detection capacity. To maximise sample purity, we also removed all sources flagged as possible contaminants in Sect. 5.2. Next, we also removed all sources flagged as multiple in Sect. 4, except for cases where the only multiplicity flag was the RUWE-based one, R (Sect. 4.4.3). This is required to preserve sources in the RUWE range that Fitton et al. (2022) claim to be populated by disc-bearing sources. Although 51% of the R labelled sample is still removed, as they were also identified as multiple systems by other methods, a modest fraction of sources with RUWE and probably real unresolved systems are left in the sample. To further minimise the influence of remaining binaries in our results, we focused our analysis on the RUWE2.5 range, which has been suggested to be the relevant range for disc-bearing sources (see Fitton et al., 2022). Our final sample was composed of 4 364 discless sources and 1 489 thick disc sources.
We employed a series of non-parametric statistical tests (Kolmogorov-Smirnov, Mann-Whitney U, and Anderson-Darling) to examine the differences between RUWE distributions of disc-bearing and discless sources, finding no support for significant differences between the two distributions. Alternatively, we have also explored other criteria for sample selection (e.g., the use of colours for thick disc selection) and other statistical approaches, such as examining disc fraction as a function of RUWE, but found no statistically significant indication of RUWE inflation due to the presence of thick discs around less evolved YSOs in our sample.
Does the variability of YSOs inflate RUWE?
Beyond extended sources, photometric variability has also been proposed as a contributor to RUWE. For example, Belokurov et al. (2020) previously uncovered a trend of inflated RUWE as a function of variability amplitude for RRLyrae and Cepheids observed with Gaia DR2. As further discussed by these authors, the inflation of RUWE by variability can be traced back to RUWE’s definition (Lindegren, 2018, eq. (4)). As its name says, RUWE is a normalised version of Gaia’s 5-parameter astrometric solution’s unit error, where a normalisation coefficient is derived from one of the modes UWE’s as a function of magnitude and colour. In variable stars where sharp changes of magnitude and colour take place, this normalisation may be incorrect, yielding spurious inflation of RUWE. Since variability is a prevalent characteristic of YSOs (Sect. 3.8), here we also address the influence of variability on Gaia’s DR3 RUWE.
We used Gaia DR3 calculated in Sect. 3.8.1 (available for of the sources in our catalogue) to address the evidence for variability influence on RUWE (e.g. Belokurov et al., 2020; Barber & Mann, 2023). As in the previous section, we minimised the incidence of binaries in the sample by removing all sources labelled as such, except for sources with only a R flag, and examining RUWE values up to 2.5. We also removed all sources flagged as possible contaminants in Sect. 5.2, which left us with 10,326 sources. Next, we bin the data at magnitude ranges (848 sources), (3,571), (3,850), and : (2,057) with the first two bins following Gaia’s window classes, and the remaining aimed at minimising the effects of the signal-to-noise variations with magnitude. For each magnitude bin, we estimate the first, second, and third rolling quartiles for the distributions of RUWE as functions of . The results are shown in Fig. 11, where variations of in RUWE are verified as a function of amplitude.
4.5 Possible Duplicates
After extensively evaluating the multiplicity of sources in our catalogue, we are now equipped to address the incidence of duplicated sources in our data compilation process. An internal match with sources’ sky positions and a 2” radius reveals that 1,864 sources are possibly duplicated. We verified that 1,687 of these had a multiplicity flag derived in this section or had their possible duplication flagged as multiple. From the remaining, 119 sources could be traced back to the publications introducing them to our dataset in Sect. 2.1, which were typically surveys including higher-resolution images, and we could confirm that multiple sources were indeed reported by their original publication, 4 had support for being independent sources from a match with Simbad, and 20 had a neighbourhood flag (Sect. 4.3) indicative of at least one neighbour within a 5” radius. Finally, only 34 sources could not be explained and are likely duplicates due to some coordinate mismatch in our historical data compilation. These sources were flagged accordingly.
5 Catalogue purity and contamination assessment
5.1 Comparison with Simbad
We used Simbad’s TAP service (on 13th December 2024) to retrieve data for 165 831 sources indexed by Simbad and located in the same field as our survey. We used a 5” searching radius to search for possible counterparts at the position of each source in our catalogue. This relatively large searching radius was chosen to account for possible offsets between coordinates adopted in the historical compilation and the ones adopted by CDS888It was noted during the data integration process (Sect. 2.1.3) that infrared source coordinates (e.g., from 2MASS) sometimes show offsets with optical coordinates (e.g., from Gaia) that cannot be explained by the different epochs of observations. This was particularly relevant for YSOs with large envelope contributions or part of multiple systems. As a cautionary example, we point to the star V 1118 Ori (NEMESIS Internal_ID 5465), a widely studied Class II YSO reported in our database with its 2MASS coordinate (2MASS J05344474-0533421). V 1118 Ori is indexed by Simbad with its ICSR J2000 coordinates from Gaia DR2, which are offset by 3.67” from its 2MASS coordinates.. We found 33 846 sources indexed by Simbad with a possible counterpart in our catalogue. 25 902 sources in our catalogue have at least one Simbad counterpart within the searching radius. All possible counterparts were recorded and are reported in our catalogue, regardless of their Simbad object type label (Tab. ), including 3 167 sources with multiple possible counterparts. We leave it to the users of the catalogue to rank possible counterparts according to their use case.
5.1.1 Young sources known as such by Simbad
We followed SIMBAD User’s Guide (v1.8) guidelines on object types to identify 19 625 sources in the OSFC field with object_type labels indicative of T Tauri Stars (TT? or TT*), Young Stellar Object (Y*?, Y*O), Ae/Be Herbig stars (Ae* or Ae?), or Orion Variables (Or*, which variable stars with irregular variability associated with eruption phenomena in young stars). We note that this list does not include objects that are arguably related to star-forming regions, such as dense cores (cor) and Herbig Haro objects (HH). Although some of these sources are indirectly included in our catalogue, as discussed in Sect. 2.1, they were considered outside the main scope of our compilation. Here, we call the first group of Simbad’s object type labels ”YSO labels”, i.e. labels that directly imply the source is a YSO. We call the second group ”star-formation labels”, i.e. labels that imply that the source is part of a star-forming region. Considering both sets of labels, 18 029 sources in our catalogue were already known by Simbad as young. Among the 7 873 sources in our catalogue with a close Sibmad counterpart but not in the two categories of labels considered, 87% were generically labelled as star (*).
5.1.2 YSOs known by Simbad but missed by our compilation
Among the YSOs known by Simbad, 2 112 were not included in our catalogue. 280 of these are located in Mon R2, which was excluded from our survey. We have retrieved the bibliographic references associated with the remaining sources and verified that 1,005 of them originate from studies with titles and abstracts focused on clumps and cores, which were considered out of our scope in Sect. 2.1. 251 sources could be traced back to CDS tables included in our compilation but were not deemed YSO candidates by the criteria in Sect. 3. The remaining sources were associated with one of the following cases. They came from all-sky variability surveys including ML-generated YSO variability flags (e.g. Samus’ et al., 2017); they came from all-sky cluster membership studies with Gaia data, whereas we have only included such surveys in our compilation when they were specifically tailored for studying the OSFC or star-forming regions (e.g. Zari et al., 2018); they came from studies focused on the OSFC but did not include tables on CDS and had too few sources for their tables to make it to the list of tables digitalised by us in Sect. 2.1.2; They came from studies focused on YSOs, but with bibliographic entries absent from the searching terms utilised in Sect. 2.1.1. Finally, we identified one bibliographic entry that adds a few dozen YSO candidates (Muench et al., 2002), which was discarded in Sect. 2.1.2 due to its focus on massive stars.
5.1.3 Sources unknown to Simbad
Our catalogue included 1 977 sources without a counterpart in Simbad. It is interesting to note that although these sources were unknown to Simbad, 90% of them had data collected with the VizieR-SED tool within a search radius of 2”. 95% of these sources had at least 4 data points in their compiled SEDs, but some had as many as 50 data points. Beyond their SEDs, 1 623 sources unknown to Simbad had measurements of at least one of the relevant observables discussed in Sect. 3, with 89% having in Sect. 3.3.1 comprising 477, 847, 140, 55, and 237 were classified as discless Class III, thin-disc Class III, Class II, Flat-Spectrum and Class 0/I respectively. The 125 sources unknown to Simbad that did not include any relevant observables related to youth other than photometry could be traced back to publications included in our compilation that selected YSOs using a CMD, with 42 of these sources from Da Rio et al. (2009), 70 from Suárez et al. (2019) and 42 from Bouy et al. (2014).
5.2 Non-YSOs and catalogue purity
In Sect. 3, we employed a positive evidence rule to compose our catalogue, where a single peer-reviewed study reporting a source as a YSO candidate was enough to grant its inclusion here. We have thus prioritised the completeness of our YSO census, rather than its purity. Consequently, YSO purity must be addressed depending on the intended use of the catalogue. Although a formal evaluation of the reliability of youth indicators of each individual source in the catalogue is beyond the scope of this paper and the subject of a follow-up study, here we discuss a series of assumptions to assess our catalogue’s degree of contamination by different types of sources. To support this discussion, Fig. 12 shows a Gaia DR3 CMD for 23,716 sources in our sample with Gaia photometry and parallaxes. As a reference, we used solar-metallicity MESA isochrones (Dotter, 2016).
5.2.1 Massive stars
We call massive stars high-mass stars that are past the PMS phase. From a combination of the adopted MESA stellar evolution models with Pecaut & Mamajek (2013) -Spectral Type scales, and considering sources at 1 Myrs, this is equivalent to sources with spectral type earlier than B3. Based on the spectral types collated in Sect. 3.1.1, 80 sources in our catalogue met this criterion (see Fig. 12 left panel), yielding a incidence of massive stars.
5.2.2 Contamination by Giants
Because background reddened giants are a known major pollutant to YSO identification studies, we wish to further add value to our catalogue by ranking possible giant contaminants. In Sect. 3.5 we collected 57 649 measurements for 12 313 sources in our catalogue. Although giant contamination in our catalogue was already hinted at in Fig. 7, the direct use of these measurements for the purpose of identifying giants is hampered by the wide spread in derivations from different surveys. Thus, we used uncertainties as a statistical weight when combining multiple literature values and restricted our analysis to 6 895 sources with either a 3 dispersion or uncertainty (if only one measurement is available) below 0.1 dex. Fig. 12 (right panel) shows the distribution of for this sample, which yields the identification of 537 likely giants with (this threshold was chosen based on distributions of values in the MESA evolutionary models with the aim of minimising overlap in distributions of giants and YSO).
To expand the giant contamination investigation to a larger portion of our catalogue, in App. E we employed giants identified based on estimation from large spectroscopic surveys to implement a Random-Forest binary classifier trained on features from Gaia DR3 to estimate the probability that sources in our catalogue have properties comparable to a known sample of Giant stars. The results of this giant classifier for our sample are shown in the middle panel of Fig. 12. We were able to derive such probabilities for 20 546 sources in our catalogue (reported in Tab. ), and 2 614 sources with high probabilities of being a Giant. We thus estimate a giant contamination level of in our catalogue. We note that among high giant probability sources, 515 sources were in the sample detected from their logg. Altogether, sources with logg indicative of giants and sources identified by our classifier correspond to 9.5% of our catalogue, suggesting a small residual incidence of unlabelled giants.
5.2.3 Extragalactic contamination
As part of another study within the framework of NEMESIS (Hernandez et al., 2024 submitted), we visually identified 11 sources with galaxy morphology in photometric images. One of these sources, 2MASS J05401945-0713593 (former Internal_ID=4281) had a counterpart in the NASA/IPAC Extragalactic Database and was removed from our compilation as soon as identified. Another source, 9334 (HOPS 292), has been pointed out as a possible Galaxy in the YSO literature. The other 9 (4330, 4331, 4547, 4608, 4609, 7319, 7471, 7492, 7533) are new identifications of Galaxies, with 8 of these previously reported as YSO by Simbad.
In App. E.2.1, we collated a sample of known extragalactic sources in the field of the OSFC, including resolved and unresolved Galaxies, AGNs and Quasi-stellar objects or Quasars (QSOs). Among this literature-based extragalactic sample, 34 sources had a close counterpart in our catalogue. The full labelled sample was used in the implementation of a Random Forest binary classifier trained to identify extragalactic sources based on photometric colours from the Gaia DR3, PanSTARRs DR2, and CatWISE surveys. The trained classifier was used to evaluate the probability of 15 855 sources in our catalogue having properties comparable to the known sample of extragalactic sources. This allowed us to identify 140 likely extragalactic sources in our catalogue, 115 of those previously unlabelled extragalactic, yielding an estimated incidence of extragalactic contaminant in our sample. Fig. 13 shows one of the parameter spaces defined by the features used in our classifier. Altogether, we identified 159 probable extragalactic sources in our sample (, suggesting a small residual incidence of extragalactic contaminants.
We note that the misclassification of extragalactic sources and YSOs happens both ways. During the collation of a labelled sample for our extragalactic classifier in App. E, we refrained from including the Gaia DR3 QSO candidates in the labelled sample due to evidence of a relevant incidence of YSOs in this sample. Following Gaia DR3’s documentation (Gaia Collaboration et al., 2023b) for selecting a high-purity QSO sample towards the OSFC, we found 81 QSO candidates in common with our YSO catalogue. Further examination of these sources on light of the diverse attributes included in our database revealed that a relevant fraction of such sources had compelling evidence for bona fide YSO classification in our catalogue, with 35 of these having spectroscopic constraints indicative of youth. Out of the 46 remaining, 30 were identified as high probability extragalactic contaminants in App. E, but 16 are unaccounted. While we opt not to include this sample in our contaminant sample, users of the catalogue who wish for a full removal of possible extragalactic sources are recommended to use the extragalactic_label and strat_label (Tab. ) to identify and further inspect such sources.
5.2.4 Further contamination
As also suggested by our comparison with Simbad, Fig. 12 also reveals contamination by low-mass MS stars. We isolated this contamination based on a 20 Myr MESA isochrone, finding 2 845 sources below the isochrone. Based on the amount of Gaia DR3 good-quality data available, this suggests a % contamination incidence.
6 Summary of Results
We have assembled the largest historical compilation of YSOs in the Orion Star Formation Complex to date, including 27 879 previously reported YSO candidates. The list of YSO candidates reported here included:
1.
Young sources identified through HRD or CMD analysis or on the bolometric luminosity-temperature relations of less evolved YSOs, including or measurements for 63% sources and Spectral Types of 41%.
2.
Sources identified in photometric and spectroscopic studies focused on infrared emission of circumstellar material, including IR classification for 35% of sources.
3.
78% of sources included spectroscopic features associated with the star-disc-interaction and accretion processes in YSOs;
4.
Youth-related spectral absorption lines such as Lithium and certain gravity-sensitive spectral lines for 22% and 48% of sources;
5.
Sources identified by kinematic studies focused on the OSFC populations including RV measurements collated for 49% of sources.
6.
Variability amplitudes for 88%, and variability periods collated for 36% of sources.
7.
X-ray emission traced back to YSOs, with X-ray luminosities and/or fluxes collated for 14% of sources.
Beyond data collated from 217 previous publications, the catalogue presented here was also complemented with panchromatic data from both photometric and spectroscopic large-scale surveys. We further added value to our catalogue by:
8.
compiling SEDs for all sources in the catalogue, with 96% of sources with at least 5 data points in their SEDs;
9.
homogeneously deriving indices and IR classes for 93% of sources;
10.
extensively evaluating the multiplicity, resulting in 43% sources in the catalogue labelled as likely multiple sources (binary candidates or visual pairs).
Due to the positive evidence approach adopted for our data collection, our catalogue prioritises completeness over purity. We estimate that of the sources in our catalogue are reliable YSOs, with 0.3% of the catalogue is composed of massive stars, 13% are likely giants contaminants, are likely extragalactic contaminants, and 12.5% are likely MS contaminants. Finally, we have estimated that the incidence of duplicates in this catalogue is below the 1% level.
As an illustrative application of the catalogue, we add to the evidence that Gaia’s RUWE parameter - commonly used for identifying unresolved binaries - may be inflated in YSOs due to their large variability.
The present catalogue is already in use by our collaborators for upcoming publications related to the NEMESIS project. (Hernandez et al., 2024 submitted) employed the catalogue to select a list of YSOs to investigate the morphology of YSOs using self-organising maps; (Marton et al., 2024b, 2025 in prep) used the catalogue as a training set for the identification of YSOs all-sky using deep learning techniques; and (Mas et al., 2025) used the YSO observables collated here to validate the results concerning the variability of YSOs observed by Gaia DR3. (Gezer et al., 2025) presented further YSO parameters by fitting YSOs’ SEDs with disc models. An ongoing study will further add value to the present catalogue through the ranking of bona fide YSOs (Roquette et al. in preparation).
Acknowledgements.
We acknowledge funding from the European Union’s Horizon 2020 research and innovation program (grant agreement No.101004141, NEMESIS). G.M. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. This research was also supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project 521 selected in 2021, Revisiting Star Formation in the Era of Big Data (https://teams.issibern.ch/starformation/) We thank Berry Holl for helpful discussions on the use of Gaia DR3 data, Lynne Hillenbrand for exchanges about large-scale collection of YSOs, and Sotiria Fotopoulou and Javier Acevedo Barroso for illuminating discussion on the identification of extragalactic sources. This work made use of Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Scikit-learn: Machine Learning in Python (Pedregosa et al., 2011), Pandas (pandas development team, 2020), Numpy (Harris et al., 2020), Natural Language Toolkit for Python (Bird et al., 2009), and Matplotlib (Hunter, 2007).References
- Abdurro’uf et al. (2022) Abdurro’uf, Accetta, K., Aerts, C., et al. 2022, ApJS, 259, 35
- Abt et al. (1991) Abt, H. A., Wang, R., & Cardona, O. 1991, ApJ, 367, 155
- Alcalá et al. (2000) Alcalá, J. M., Covino, E., Torres, G., et al. 2000, A&A, 353, 186
- Alcalá et al. (2017) Alcalá, J. M., Manara, C. F., Natta, A., et al. 2017, A&A, 600, A20
- Alcalá et al. (2014) Alcalá, J. M., Natta, A., Manara, C. F., et al. 2014, A&A, 561, A2
- Alcalá et al. (2004) Alcalá, J. M., Wachter, S., Covino, E., et al. 2004, A&A, 416, 677
- Alecian et al. (2009) Alecian, E., Wade, G. A., Catala, C., et al. 2009, MNRAS, 400, 354
- Ali & Noriega-Crespo (2004) Ali, B. & Noriega-Crespo, A. 2004, ApJ, 613, 374
- Andrae et al. (2023) Andrae, R., Fouesneau, M., Sordo, R., et al. 2023, A&A, 674, A27
- Andre et al. (1993) Andre, P., Ward-Thompson, D., & Barsony, M. 1993, ApJ, 406, 122
- Aristarán et al. (v. 1.2.1) Aristarán, M., Tigas, M., B., M. J., Das, J., & Frackman, D. v. 1.2.1, Tabula
- Arun et al. (2021) Arun, R., Mathew, B., Rengaswamy, S., et al. 2021, MNRAS, 501, 1243
- Astropy Collaboration et al. (2022) Astropy Collaboration, Price-Whelan, A. M., Lim, P. L., et al. 2022, ApJ, 935, 167
- Astropy Collaboration et al. (2018) Astropy Collaboration, Price-Whelan, A. M., Sipőcz, B. M., et al. 2018, AJ, 156, 123
- Astropy Collaboration et al. (2013) Astropy Collaboration, Robitaille, T. P., Tollerud, E. J., et al. 2013, A&A, 558, A33
- Azevedo et al. (2006) Azevedo, R., Calvet, N., Hartmann, L., et al. 2006, A&A, 456, 225
- Baines et al. (2004) Baines, D., Oudmaijer, R. D., Mora, A., et al. 2004, MNRAS, 353, 697
- Baldovin-Saavedra et al. (2009) Baldovin-Saavedra, C., Audard, M., Duchêne, G., et al. 2009, ApJ, 697, 493
- Bally (2008) Bally, J. 2008, in Handbook of Star Forming Regions, Volume I, ed. B. Reipurth, Vol. 4, 459
- Baraffe & Chabrier (2010) Baraffe, I. & Chabrier, G. 2010, A&A, 521, A44
- Barber & Mann (2023) Barber, M. G. & Mann, A. W. 2023, ApJ, 953, 127
- Barrado et al. (2011) Barrado, D., Stelzer, B., Morales-Calderón, M., et al. 2011, A&A, 526, A21
- Barrado Y Navascués et al. (2007) Barrado Y Navascués, D., Bayo, A., Morales-Calderón, M., et al. 2007, A&A, 468, L5
- Barrado y Navascués & Martín (2003) Barrado y Navascués, D. & Martín, E. L. 2003, AJ, 126, 2997
- Barrado y Navascués et al. (2004) Barrado y Navascués, D., Stauffer, J. R., Bouvier, J., Jayawardhana, R., & Cuillandre, J.-C. 2004, ApJ, 610, 1064
- Barrado y Navascués et al. (2007) Barrado y Navascués, D., Stauffer, J. R., Morales-Calderón, M., et al. 2007, ApJ, 664, 481
- Basri & Batalha (1990) Basri, G. & Batalha, C. 1990, ApJ, 363, 654
- Basri & Bertout (1989) Basri, G. & Bertout, C. 1989, ApJ, 341, 340
- Bayo et al. (2012) Bayo, A., Barrado, D., Huélamo, N., et al. 2012, A&A, 547, A80
- Bayo et al. (2011) Bayo, A., Barrado, D., Stauffer, J., et al. 2011, A&A, 536, A63
- Bayo et al. (2008) Bayo, A., Rodrigo, C., Barrado Y Navascués, D., et al. 2008, A&A, 492, 277
- Béjar et al. (2001) Béjar, V. J. S., Martín, E. L., Zapatero Osorio, M. R., et al. 2001, ApJ, 556, 830
- Béjar et al. (2011) Béjar, V. J. S., Zapatero Osorio, M. R., Rebolo, R., et al. 2011, ApJ, 743, 64
- Belokurov et al. (2020) Belokurov, V., Penoyre, Z., Oh, S., et al. 2020, MNRAS, 496, 1922
- Bianchi et al. (2017) Bianchi, L., Shiao, B., & Thilker, D. 2017, ApJS, 230, 24
- Biazzo et al. (2009) Biazzo, K., Melo, C. H. F., Pasquini, L., et al. 2009, A&A, 508, 1301
- Bird et al. (2009) Bird, S., Klein, E., & Loper, E. 2009, Natural Language Processing with Python: Analyzing Text with the Natural Language Toolkit (O’Reilly Media)
- Birko et al. (2019) Birko, D., Zwitter, T., Grebel, E. K., et al. 2019, AJ, 158, 155
- Biteau (2021) Biteau, J. 2021, ApJS, 256, 15
- Bolton et al. (1998) Bolton, C. T., Harmanec, P., Lyons, R. W., Odell, A. P., & Pyper, D. M. 1998, A&A, 337, 183
- Bossi et al. (1989) Bossi, M., Gaspani, A., Scardia, M., & Tadini, M. 1989, A&A, 222, 117
- Bouma et al. (2020) Bouma, L. G., Winn, J. N., Ricker, G. R., et al. 2020, AJ, 160, 86
- Bouvier et al. (2007) Bouvier, J., Alencar, S. H. P., Harries, T. J., Johns-Krull, C. M., & Romanova, M. M. 2007, in Protostars and Planets V, ed. B. Reipurth, D. Jewitt, & K. Keil, 479
- Bouy et al. (2014) Bouy, H., Alves, J., Bertin, E., Sarro, L. M., & Barrado, D. 2014, A&A, 564, A29
- Bradley et al. (2024) Bradley, L., Sipőcz, B., Robitaille, T., et al. 2024, astropy/photutils: 1.13.0
- Breiman (2001) Breiman, L. 2001, Machine Learning, 45, 5
- Briceño et al. (2005) Briceño, C., Calvet, N., Hernández, J., et al. 2005, AJ, 129, 907
- Briceño et al. (2019) Briceño, C., Calvet, N., Hernández, J., et al. 2019, AJ, 157, 85
- Briceño et al. (2007) Briceño, C., Hartmann, L., Hernández, J., et al. 2007, ApJ, 661, 1119
- Broos et al. (2013) Broos, P. S., Getman, K. V., Povich, M. S., et al. 2013, ApJS, 209, 32
- Bruursema et al. (2023) Bruursema, J., Vrba, F., Munn, J., et al. 2023, in American Astronomical Society Meeting Abstracts, Vol. 242, American Astronomical Society Meeting Abstracts, 118.08
- Buder et al. (2024) Buder, S., Kos, J., Wang, E. X., et al. 2024, arXiv e-prints, arXiv:2409.19858
- Burningham et al. (2005) Burningham, B., Naylor, T., Littlefair, S. P., & Jeffries, R. D. 2005, MNRAS, 356, 1583
- Caballero (2007) Caballero, J. A. 2007, A&A, 466, 917
- Caballero (2008) Caballero, J. A. 2008, A&A, 478, 667
- Caballero et al. (2012) Caballero, J. A., Cabrera-Lavers, A., García-Álvarez, D., & Pascual, S. 2012, A&A, 546, A59
- Caballero et al. (2019) Caballero, J. A., de Burgos, A., Alonso-Floriano, F. J., et al. 2019, A&A, 629, A114
- Caballero et al. (2018) Caballero, J. A., Novalbos, I., Tobal, T., & Miret, F. X. 2018, Astronomische Nachrichten, 339, 60
- Caballero & Solano (2008) Caballero, J. A. & Solano, E. 2008, A&A, 485, 931
- Cao et al. (2022) Cao, L., Pinsonneault, M. H., Hillenbrand, L. A., & Kuhn, M. A. 2022, ApJ, 924, 84
- Cargile et al. (2008) Cargile, P. A., Stassun, K. G., & Mathieu, R. D. 2008, ApJ, 674, 329
- Carnerero et al. (2023) Carnerero, M. I., Raiteri, C. M., Rimoldini, L., et al. 2023, A&A, 674, A24
- Carpenter et al. (2001) Carpenter, J. M., Hillenbrand, L. A., & Skrutskie, M. F. 2001, AJ, 121, 3160
- Casey et al. (2017) Casey, A., Sudilovsky, V., Barentsen, G., et al. 2017, A Python Module to Interact with NASA’s ADS that Doesn’t SuckTM, https://github.com/andycasey/ads
- Castro-Ginard et al. (2024) Castro-Ginard, A., Penoyre, Z., Casey, A. R., et al. 2024, A&A, 688, A1
- Chambers et al. (2016) Chambers, K. C., Magnier, E. A., Metcalfe, N., et al. 2016, arXiv e-prints, arXiv:1612.05560
- Chen et al. (1995) Chen, H., Myers, P. C., Ladd, E. F., & Wood, D. O. S. 1995, ApJ, 445, 377
- Cody & Hillenbrand (2010) Cody, A. M. & Hillenbrand, L. A. 2010, ApJS, 191, 389
- Cody et al. (2014) Cody, A. M., Stauffer, J., Baglin, A., et al. 2014, AJ, 147, 82
- Cohen et al. (2003) Cohen, M., Wheaton, W. A., & Megeath, S. T. 2003, AJ, 126, 1090
- Connelley & Greene (2010) Connelley, M. S. & Greene, T. P. 2010, AJ, 140, 1214
- Cook et al. (2024) Cook, T. L., Bandi, B., Philipsborn, S., et al. 2024, MNRAS, 535, 2129
- Correia et al. (2013) Correia, S., Duchêne, G., Reipurth, B., et al. 2013, A&A, 557, A63
- Cottle et al. (2018) Cottle, J. N., Covey, K. R., Suárez, G., et al. 2018, ApJS, 236, 27
- Covino et al. (2000) Covino, E., Catalano, S., Frasca, A., et al. 2000, A&A, 361, L49
- Covino et al. (2001) Covino, E., Melo, C., Alcalá, J. M., et al. 2001, A&A, 375, 130
- Creevey et al. (2023) Creevey, O. L., Sordo, R., Pailler, F., et al. 2023, A&A, 674, A26
- Da Rio et al. (2012) Da Rio, N., Robberto, M., Hillenbrand, L. A., Henning, T., & Stassun, K. G. 2012, ApJ, 748, 14
- Da Rio et al. (2009) Da Rio, N., Robberto, M., Soderblom, D. R., et al. 2009, ApJS, 183, 261
- Da Rio et al. (2010) Da Rio, N., Robberto, M., Soderblom, D. R., et al. 2010, ApJ, 722, 1092
- Da Rio et al. (2016) Da Rio, N., Tan, J. C., Covey, K. R., et al. 2016, ApJ, 818, 59
- Daemgen et al. (2012) Daemgen, S., Correia, S., & Petr-Gotzens, M. G. 2012, A&A, 540, A46
- Damian et al. (2023) Damian, B., Jose, J., Biller, B., et al. 2023, ApJ, 951, 139
- Davies et al. (2018) Davies, C. L., Kreplin, A., Kluska, J., Hone, E., & Kraus, S. 2018, MNRAS, 474, 5406
- De Angeli et al. (2023) De Angeli, F., Weiler, M., Montegriffo, P., et al. 2023, A&A, 674, A2
- De Furio et al. (2022a) De Furio, M., Liu, C., Meyer, M. R., et al. 2022a, ApJ, 941, 161
- De Furio et al. (2022b) De Furio, M., Meyer, M. R., Reiter, M., et al. 2022b, ApJ, 925, 112
- De Furio et al. (2019) De Furio, M., Reiter, M., Meyer, M. R., et al. 2019, ApJ, 886, 95
- Delchambre et al. (2023) Delchambre, L., Bailer-Jones, C. A. L., Bellas-Velidis, I., et al. 2023, A&A, 674, A31
- Distefano et al. (2023) Distefano, E., Lanzafame, A. C., Brugaletta, E., et al. 2023, A&A, 674, A20
- Dolan & Mathieu (2001) Dolan, C. J. & Mathieu, R. D. 2001, AJ, 121, 2124
- Dotter (2016) Dotter, A. 2016, ApJS, 222, 8
- Dowler et al. (2019) Dowler, P., Rixon, G., Tody, D., & Demleitner, M. 2019, Table Access Protocol Version 1.1, IVOA Recommendation 27 September 2019
- Downes et al. (2014) Downes, J. J., Briceño, C., Mateu, C., et al. 2014, MNRAS, 444, 1793
- Downes et al. (2015) Downes, J. J., Román-Zúñiga, C., Ballesteros-Paredes, J., et al. 2015, MNRAS, 450, 3490
- Duchêne et al. (2018) Duchêne, G., Lacour, S., Moraux, E., Goodwin, S., & Bouvier, J. 2018, MNRAS, 478, 1825
- Ducourant et al. (2023) Ducourant, C., Krone-Martins, A., Galluccio, L., et al. 2023, A&A, 674, A11
- Dunham et al. (2014) Dunham, M. M., Stutz, A. M., Allen, L. E., et al. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 195–218
- Dupree et al. (1992) Dupree, A. K., Sasselov, D. D., & Lester, J. B. 1992, ApJ, 387, L85
- Dye et al. (2018) Dye, S., Lawrence, A., Read, M. A., et al. 2018, MNRAS, 473, 5113
- El-Badry et al. (2021) El-Badry, K., Rix, H.-W., & Heintz, T. M. 2021, MNRAS, 506, 2269
- Espaillat et al. (2014) Espaillat, C., Muzerolle, J., Najita, J., et al. 2014, in Protostars and Planets VI, ed. H. Beuther, R. S. Klessen, C. P. Dullemond, & T. Henning, 497–520
- Fabricius et al. (2021) Fabricius, C., Luri, X., Arenou, F., et al. 2021, A&A, 649, A5
- Fang et al. (2021) Fang, M., Kim, J. S., Pascucci, I., & Apai, D. 2021, ApJ, 908, 49
- Fang et al. (2017) Fang, M., Kim, J. S., Pascucci, I., et al. 2017, AJ, 153, 188
- Fang et al. (2013) Fang, M., Kim, J. S., van Boekel, R., et al. 2013, ApJS, 207, 5
- Fang et al. (2014) Fang, M., Sicilia-Aguilar, A., Roccatagliata, V., et al. 2014, A&A, 570, A118
- Fang et al. (2009) Fang, M., van Boekel, R., Wang, W., et al. 2009, A&A, 504, 461
- Fazio et al. (2004) Fazio, G. G., Hora, J. L., Allen, L. E., et al. 2004, ApJS, 154, 10
- Federman et al. (2023) Federman, S., Megeath, S. T., Tobin, J. J., et al. 2023, ApJ, 944, 49
- Feigelson et al. (2002) Feigelson, E. D., Broos, P., Gaffney, III, J. A., et al. 2002, ApJ, 574, 258
- Feigelson & Montmerle (1999) Feigelson, E. D. & Montmerle, T. 1999, ARA&A, 37, 363
- Fernandez et al. (2017) Fernandez, M. A., Covey, K. R., De Lee, N., et al. 2017, PASP, 129, 084201
- Fűrész et al. (2008) Fűrész, G., Hartmann, L. W., Megeath, S. T., Szentgyorgyi, A. H., & Hamden, E. T. 2008, ApJ, 676, 1109
- Fischer et al. (2020) Fischer, W. J., Megeath, S. T., Furlan, E., et al. 2020, ApJ, 905, 119
- Fitton et al. (2022) Fitton, S., Tofflemire, B. M., & Kraus, A. L. 2022, Research Notes of the American Astronomical Society, 6, 18
- Flaccomio et al. (2003) Flaccomio, E., Damiani, F., Micela, G., et al. 2003, ApJ, 582, 382
- Flaischlen et al. (2022) Flaischlen, S., Preibisch, T., Kluge, M., Manara, C. F., & Ercolano, B. 2022, A&A, 666, A55
- Flesch (2023) Flesch, E. W. 2023, The Open Journal of Astrophysics, 6, 49
- Flewelling et al. (2020) Flewelling, H. A., Magnier, E. A., Chambers, K. C., et al. 2020, ApJS, 251, 7
- Folha & Emerson (1999) Folha, D. F. M. & Emerson, J. P. 1999, A&A, 352, 517
- Franciosini et al. (2006) Franciosini, E., Pallavicini, R., & Sanz-Forcada, J. 2006, A&A, 446, 501
- Franciosini & Sacco (2011) Franciosini, E. & Sacco, G. G. 2011, A&A, 530, A150
- Franciosini et al. (2022) Franciosini, E., Tognelli, E., Degl’Innocenti, S., et al. 2022, A&A, 659, A85
- Frasca et al. (2021) Frasca, A., Boffin, H. M. J., Manara, C. F., et al. 2021, A&A, 656, A138
- Frasca et al. (2009) Frasca, A., Covino, E., Spezzi, L., et al. 2009, A&A, 508, 1313
- Furlan et al. (2016) Furlan, E., Fischer, W. J., Ali, B., et al. 2016, ApJS, 224, 5
- Gagne & Caillault (1994) Gagne, M. & Caillault, J.-P. 1994, ApJ, 437, 361
- Gaia Collaboration et al. (2023a) Gaia Collaboration, Arenou, F., Babusiaux, C., et al. 2023a, A&A, 674, A34
- Gaia Collaboration et al. (2023b) Gaia Collaboration, Bailer-Jones, C. A. L., Teyssier, D., et al. 2023b, A&A, 674, A41
- (131) Gaia ESA Archive. , official webpage, https://gea.esac.esa.int/archive/
- Genova et al. (2000) Genova, F., Egret, D., Bienaymé, O., et al. 2000, A&AS, 143, 1
- Getman et al. (2017) Getman, K. V., Broos, P. S., Kuhn, M. A., et al. 2017, ApJS, 229, 28
- Getman et al. (2005a) Getman, K. V., Feigelson, E. D., Grosso, N., et al. 2005a, ApJS, 160, 353
- Getman et al. (2005b) Getman, K. V., Flaccomio, E., Broos, P. S., et al. 2005b, ApJS, 160, 319
- Getman et al. (2018) Getman, K. V., Kuhn, M. A., Feigelson, E. D., et al. 2018, MNRAS, 477, 298
- Gezer et al. (2025) Gezer, I., Marton, G., Roquette, J., et al. 2025, A&A, 696, A196
- Gies et al. (1996) Gies, D. R., Barry, D. J., Bagnuolo, Jr., W. G., Sowers, J., & Thaller, M. L. 1996, ApJ, 469, 884
- Gilmore et al. (2022) Gilmore, G., Randich, S., Worley, C. C., et al. 2022, A&A, 666, A120
- Ginsburg et al. (2019) Ginsburg, A., Sipőcz, B. M., Brasseur, C. E., et al. 2019, AJ, 157, 98
- Gómez Maqueo Chew et al. (2012) Gómez Maqueo Chew, Y., Stassun, K. G., Prša, A., et al. 2012, ApJ, 745, 58
- Gosset et al. (2024) Gosset, E., Damerdji, Y., Morel, T., et al. 2024, arXiv e-prints, arXiv:2410.14372
- Grant et al. (2018) Grant, S. L., Espaillat, C. C., Megeath, S. T., et al. 2018, ApJ, 863, 13
- GRAVITY Collaboration et al. (2018) GRAVITY Collaboration, Karl, M., Pfuhl, O., et al. 2018, A&A, 620, A116
- Green et al. (2023) Green, M. J., Maoz, D., Mazeh, T., et al. 2023, MNRAS, 522, 29
- Greene & Lada (1996) Greene, T. P. & Lada, C. J. 1996, ApJ, 461, 345
- Greene et al. (1994) Greene, T. P., Wilking, B. A., Andre, P., Young, E. T., & Lada, C. J. 1994, ApJ, 434, 614
- Grinin et al. (2010) Grinin, V. P., Rostopchina, A. N., Barsunova, O. Y., & Demidova, T. V. 2010, Astrophysics, 53, 367
- Großschedl et al. (2019) Großschedl, J. E., Alves, J., Teixeira, P. S., et al. 2019, A&A, 622, A149
- Gutermuth et al. (2009) Gutermuth, R. A., Megeath, S. T., Myers, P. C., et al. 2009, ApJS, 184, 18
- Habel et al. (2021) Habel, N. M., Megeath, S. T., Booker, J. J., et al. 2021, ApJ, 911, 153
- Halbwachs et al. (2023) Halbwachs, J.-L., Pourbaix, D., Arenou, F., et al. 2023, A&A, 674, A9
- Hamann (1994) Hamann, F. 1994, ApJS, 93, 485
- Hamann & Persson (1992) Hamann, F. & Persson, S. E. 1992, ApJS, 82, 247
- Harris et al. (2020) Harris, C. R., Millman, K. J., van der Walt, S. J., et al. 2020, Nature, 585, 357
- Hartigan et al. (1995) Hartigan, P., Edwards, S., & Ghandour, L. 1995, ApJ, 452, 736
- Hartmann et al. (1994) Hartmann, L., Hewett, R., & Calvet, N. 1994, ApJ, 426, 669
- Hasenberger et al. (2016) Hasenberger, B., Forbrich, J., Alves, J., et al. 2016, A&A, 593, A7
- Herbig (1962) Herbig, G. H. 1962, Advances in Astronomy and Astrophysics, 1, 47
- Herbig & Griffin (2006) Herbig, G. H. & Griffin, R. F. 2006, AJ, 132, 1763
- Herbst et al. (2002) Herbst, W., Bailer-Jones, C. A. L., Mundt, R., Meisenheimer, K., & Wackermann, R. 2002, A&A, 396, 513
- Herbst et al. (1994) Herbst, W., Herbst, D. K., Grossman, E. J., & Weinstein, D. 1994, AJ, 108, 1906
- Herbst et al. (2000) Herbst, W., Rhode, K. L., Hillenbrand, L. A., & Curran, G. 2000, AJ, 119, 261
- Hernandez et al. (2024 submitted) Hernandez, D., Dionatos, O., Audard, M., et al. 2024 submitted, A&A
- Hernández et al. (2007a) Hernández, J., Calvet, N., Briceño, C., et al. 2007a, ApJ, 671, 1784
- Hernández et al. (2009) Hernández, J., Calvet, N., Hartmann, L., et al. 2009, ApJ, 707, 705
- Hernández et al. (2014) Hernández, J., Calvet, N., Perez, A., et al. 2014, ApJ, 794, 36
- Hernández et al. (2007b) Hernández, J., Hartmann, L., Megeath, T., et al. 2007b, ApJ, 662, 1067
- Hernández et al. (2010) Hernández, J., Morales-Calderon, M., Calvet, N., et al. 2010, ApJ, 722, 1226
- Hernández et al. (2023) Hernández, J., Zamudio, L. F., Briceño, C., et al. 2023, AJ, 165, 205
- Hesser et al. (1976) Hesser, J. E., Walborn, N. R., & Ugarte, P. P. 1976, Nature, 262, 116
- Hillenbrand (1997) Hillenbrand, L. A. 1997, AJ, 113, 1733
- Hillenbrand et al. (2013) Hillenbrand, L. A., Hoffer, A. S., & Herczeg, G. J. 2013, AJ, 146, 85
- Hillenbrand et al. (1998) Hillenbrand, L. A., Strom, S. E., Calvet, N., et al. 1998, AJ, 116, 1816
- Holl et al. (2023a) Holl, B., Fabricius, C., Portell, J., et al. 2023a, A&A, 674, A25
- Holl et al. (2023b) Holl, B., Sozzetti, A., Sahlmann, J., et al. 2023b, A&A, 674, A10
- Hourihane et al. (2023) Hourihane, A., François, P., Worley, C. C., et al. 2023, A&A, 676, A129
- Hsu et al. (2012) Hsu, W.-H., Hartmann, L., Allen, L., et al. 2012, ApJ, 752, 59
- Hsu et al. (2013) Hsu, W.-H., Hartmann, L., Allen, L., et al. 2013, ApJ, 764, 114
- Hunt & Reffert (2021) Hunt, E. L. & Reffert, S. 2021, A&A, 646, A104
- Hunter (2007) Hunter, J. D. 2007, Computing in Science & Engineering, 9, 90
- Ingleby et al. (2011) Ingleby, L., Calvet, N., Hernández, J., et al. 2011, AJ, 141, 127
- Ingraham et al. (2014) Ingraham, P., Albert, L., Doyon, R., & Artigau, E. 2014, ApJ, 782, 8
- (184) IRSA. , NASA/IPAC Infrared Science Archive, https://irsa.ipac.caltech.edu/frontpage/
- (185) IRSA WISE Image Service. , Wide-field Infrared Survey Explorer (WISE) and NEOWISE, https://irsa.ipac.caltech.edu/Missions/wise.html
- Irwin et al. (2007) Irwin, J., Aigrain, S., Hodgkin, S., et al. 2007, MNRAS, 380, 541
- Jack (2019) Jack, D. 2019, Astronomische Nachrichten, 340, 386
- Jackson et al. (2020) Jackson, R. J., Jeffries, R. D., Wright, N. J., et al. 2020, MNRAS, 496, 4701
- Jaehnig et al. (2017) Jaehnig, K., Bird, J. C., Stassun, K. G., et al. 2017, ApJ, 851, 14
- Jayasinghe et al. (2018) Jayasinghe, T., Kochanek, C. S., Stanek, K. Z., et al. 2018, MNRAS, 477, 3145
- Jeffries (2014) Jeffries, R. D. 2014, in EAS Publications Series, Vol. 65, EAS Publications Series, ed. Y. Lebreton, D. Valls-Gabaud, & C. Charbonnel, 289–325
- Jerabkova et al. (2019) Jerabkova, T., Beccari, G., Boffin, H. M. J., et al. 2019, A&A, 627, A57
- Joy (1945) Joy, A. H. 1945, ApJ, 102, 168
- Karim et al. (2016) Karim, T., Stassun, K. G., Briceño, C., et al. 2016, AJ, 152, 198
- Kenyon et al. (2005) Kenyon, M. J., Jeffries, R. D., Naylor, T., Oliveira, J. M., & Maxted, P. F. L. 2005, MNRAS, 356, 89
- Kharchenko (2001) Kharchenko, N. V. 2001, Kinematika i Fizika Nebesnykh Tel, 17, 409
- Kim et al. (2019) Kim, D., Lu, J. R., Konopacky, Q., et al. 2019, AJ, 157, 109
- Kim et al. (2016) Kim, K. H., Watson, D. M., Manoj, P., et al. 2016, ApJS, 226, 8
- Kim et al. (2013) Kim, K. H., Watson, D. M., Manoj, P., et al. 2013, ApJ, 769, 149
- Koenig et al. (2015) Koenig, X., Hillenbrand, L. A., Padgett, D. L., & DeFelippis, D. 2015, AJ, 150, 100
- Koenig & Leisawitz (2014) Koenig, X. P. & Leisawitz, D. T. 2014, ApJ, 791, 131
- Köhler et al. (2006) Köhler, R., Petr-Gotzens, M. G., McCaughrean, M. J., et al. 2006, A&A, 458, 461
- Kos et al. (2021) Kos, J., Bland-Hawthorn, J., Buder, S., et al. 2021, MNRAS, 506, 4232
- Kounkel et al. (2019) Kounkel, M., Covey, K., Moe, M., et al. 2019, AJ, 157, 196
- Kounkel et al. (2020) Kounkel, M., Covey, K., & Stassun, K. G. 2020, AJ, 160, 279
- Kounkel et al. (2018) Kounkel, M., Covey, K., Suárez, G., et al. 2018, AJ, 156, 84
- Kounkel et al. (2017a) Kounkel, M., Hartmann, L., Calvet, N., & Megeath, T. 2017a, AJ, 154, 29
- Kounkel et al. (2017b) Kounkel, M., Hartmann, L., Loinard, L., et al. 2017b, ApJ, 834, 142
- Kounkel et al. (2017c) Kounkel, M., Hartmann, L., Mateo, M., & Bailey, John I., I. 2017c, ApJ, 844, 138
- Kounkel et al. (2016a) Kounkel, M., Hartmann, L., Tobin, J. J., et al. 2016a, ApJ, 821, 8
- Kounkel et al. (2016b) Kounkel, M., Megeath, S. T., Poteet, C. A., Fischer, W. J., & Hartmann, L. 2016b, ApJ, 821, 52
- Kounkel et al. (2022) Kounkel, M., Stassun, K. G., Bouma, L. G., et al. 2022, AJ, 164, 137
- Kovalev et al. (2024) Kovalev, M., Zhou, Z., Chen, X., & Han, Z. 2024, MNRAS, 527, 521
- Kozlova et al. (1995) Kozlova, O. V., Grinin, V. P., & Rostopchina, A. N. 1995, Astronomical and Astrophysical Transactions, 8, 249
- Kryukova et al. (2012) Kryukova, E., Megeath, S. T., Gutermuth, R. A., et al. 2012, AJ, 144, 31
- Kuhn et al. (2014) Kuhn, M. A., Feigelson, E. D., Getman, K. V., et al. 2014, ApJ, 787, 107
- Kuhn et al. (2013) Kuhn, M. A., Povich, M. S., Luhman, K. L., et al. 2013, ApJS, 209, 29
- Kurosawa et al. (2011) Kurosawa, R., Romanova, M. M., & Harries, T. J. 2011, MNRAS, 416, 2623
- Kurtz et al. (2000) Kurtz, M. J., Eichhorn, G., Accomazzi, A., et al. 2000, A&AS, 143, 41
- Lada (1987) Lada, C. J. 1987, in Star Forming Regions, ed. M. Peimbert & J. Jugaku, Vol. 115, 1
- (221) LAMOST DR10 official webpage. , v. 2.0, https://www.lamost.org/dr10/v2.0/
- Lanzafame et al. (2023) Lanzafame, A. C., Brugaletta, E., Frémat, Y., et al. 2023, A&A, 674, A30
- Laos et al. (2021) Laos, S., Greene, T. P., Najita, J. R., & Stassun, K. G. 2021, ApJ, 921, 110
- Lavail et al. (2020) Lavail, A., Kochukhov, O., Hussain, G. A. J., et al. 2020, MNRAS, 497, 632
- Lawrence et al. (2007) Lawrence, A., Warren, S. J., Almaini, O., et al. 2007, MNRAS, 379, 1599
- Lee & Chen (2007) Lee, H.-T. & Chen, W. P. 2007, ApJ, 657, 884
- Leone et al. (2010) Leone, F., Bohlender, D. A., Bolton, C. T., et al. 2010, MNRAS, 401, 2739
- Lewis & Lada (2016) Lewis, J. A. & Lada, C. J. 2016, ApJ, 825, 91
- Lima et al. (2010) Lima, G. H. R. A., Alencar, S. H. P., Calvet, N., Hartmann, L., & Muzerolle, J. 2010, A&A, 522, A104
- Lindegren (2018) Lindegren, L. 2018
- Lindegren et al. (2021) Lindegren, L., Klioner, S. A., Hernández, J., et al. 2021, A&A, 649, A2
- Lloyd & Stickland (1999) Lloyd, C. & Stickland, D. J. 1999, Information Bulletin on Variable Stars, 4809, 1
- Logan & Fotopoulou (2020) Logan, C. H. A. & Fotopoulou, S. 2020, A&A, 633, A154
- López-García et al. (2013) López-García, M. A., López-Santiago, J., Albacete-Colombo, J. F., Pérez-González, P. G., & de Castro, E. 2013, MNRAS, 429, 775
- López-Santiago & Caballero (2008) López-Santiago, J. & Caballero, J. A. 2008, A&A, 491, 961
- Lourens et al. (2024) Lourens, M. A. A., Trager, S. C., Kim, Y., Telea, A. C., & Roerdink, J. B. T. M. 2024, A&A, 690, A224
- Luhman et al. (2008a) Luhman, K. L., Allen, L. E., Allen, P. R., et al. 2008a, ApJ, 675, 1375
- Luhman et al. (2008b) Luhman, K. L., Hernández, J., Downes, J. J., Hartmann, L., & Briceño, C. 2008b, ApJ, 688, 362
- Mace et al. (2009) Mace, G. N., Prato, L., Wasserman, L. H., et al. 2009, AJ, 137, 3487
- Madarász et al. (2025) Madarász, M., Marton, G., Gezer, I., et al. 2025, A&A, 696, A37
- Maheswar et al. (2003) Maheswar, G., Manoj, P., & Bhatt, H. C. 2003, A&A, 402, 963
- Mainzer et al. (2014) Mainzer, A., Bauer, J., Cutri, R. M., et al. 2014, ApJ, 792, 30
- Mairs et al. (2019) Mairs, S., Lalchand, B., Bower, G. C., et al. 2019, ApJ, 871, 72
- Makarov et al. (2014) Makarov, D., Prugniel, P., Terekhova, N., Courtois, H., & Vauglin, I. 2014, A&A, 570, A13
- Malkov et al. (2016) Malkov, O., Karchevsky, A., Kaygorodov, P., & Kovaleva, D. 2016, Baltic Astronomy, 25, 49
- Manara et al. (2012) Manara, C. F., Robberto, M., Da Rio, N., et al. 2012, ApJ, 755, 154
- Manset & Bastien (2002) Manset, N. & Bastien, P. 2002, AJ, 124, 1089
- Manzo-Martínez et al. (2020) Manzo-Martínez, E., Calvet, N., Hernández, J., et al. 2020, ApJ, 893, 56
- Marocco et al. (2021) Marocco, F., Eisenhardt, P. R. M., Fowler, J. W., et al. 2021, ApJS, 253, 8
- Marschall & Mathieu (1988) Marschall, L. A. & Mathieu, R. D. 1988, AJ, 96, 1956
- Martin et al. (1998) Martin, E. L., Montmerle, T., Gregorio-Hetem, J., & Casanova, S. 1998, MNRAS, 300, 733
- Marton et al. (2019) Marton, G., Ábrahám, P., Szegedi-Elek, E., et al. 2019, MNRAS, 487, 2522
- Marton et al. (2024a) Marton, G., Gezer, I., Madarász, M., et al. 2024a, A&A, 688, A203
- Marton et al. (2025 in prep) Marton, G., Roquette, J., Gezer, I., et al. 2025 in prep, A&A
- Marton et al. (2024b) Marton, G., Roquette, J., Madarász, M., et al. 2024b, in EAS2024, European Astronomical Society Annual Meeting, 1922
- Marton et al. (2016) Marton, G., Tóth, L. V., Paladini, R., et al. 2016, MNRAS, 458, 3479
- Mas et al. (2025) Mas, C., Roquette, J., Audard, M., et al. 2025, arXiv e-prints, arXiv:2503.23530
- Mathieu et al. (1991) Mathieu, R. D., Adams, F. C., & Latham, D. W. 1991, AJ, 101, 2184
- Maxted et al. (2008) Maxted, P. F. L., Jeffries, R. D., Oliveira, J. M., Naylor, T., & Jackson, R. J. 2008, MNRAS, 385, 2210
- McBride & Kounkel (2019) McBride, A. & Kounkel, M. 2019, ApJ, 884, 6
- McClure et al. (2010) McClure, M. K., Furlan, E., Manoj, P., et al. 2010, ApJS, 188, 75
- McMahon et al. (2013) McMahon, R. G., Banerji, M., Gonzalez, E., et al. 2013, The Messenger, 154, 35
- Megeath et al. (2012) Megeath, S. T., Gutermuth, R., Muzerolle, J., et al. 2012, AJ, 144, 192
- Megeath et al. (2016) Megeath, S. T., Gutermuth, R., Muzerolle, J., et al. 2016, AJ, 151, 5
- Meingast et al. (2016) Meingast, S., Alves, J., Mardones, D., et al. 2016, A&A, 587, A153
- Messina et al. (2016) Messina, S., Parihar, P., Biazzo, K., et al. 2016, MNRAS, 457, 3372
- Messina et al. (2017) Messina, S., Parihar, P., & Distefano, E. 2017, MNRAS, 468, 931
- Mohanty et al. (2009) Mohanty, S., Stassun, K. G., & Mathieu, R. D. 2009, ApJ, 697, 713
- Mookerjea & Sandell (2009) Mookerjea, B. & Sandell, G. 2009, ApJ, 706, 896
- Morales-Calderón et al. (2011) Morales-Calderón, M., Stauffer, J. R., Hillenbrand, L. A., et al. 2011, ApJ, 733, 50
- Morales-Calderón et al. (2012) Morales-Calderón, M., Stauffer, J. R., Stassun, K. G., et al. 2012, ApJ, 753, 149
- Mowlavi et al. (2023) Mowlavi, N., Holl, B., Lecoeur-Taïbi, I., et al. 2023, A&A, 674, A16
- Mowlavi et al. (2021) Mowlavi, N., Rimoldini, L., Evans, D. W., et al. 2021, A&A, 648, A44
- Muench et al. (2008) Muench, A., Getman, K., Hillenbrand, L., & Preibisch, T. 2008, in Handbook of Star Forming Regions, Volume I, ed. B. Reipurth, Vol. 4, 483
- Muench et al. (2002) Muench, A. A., Lada, E. A., Lada, C. J., & Alves, J. 2002, ApJ, 573, 366
- Muzerolle et al. (1998a) Muzerolle, J., Calvet, N., & Hartmann, L. 1998a, ApJ, 492, 743
- Muzerolle et al. (2001) Muzerolle, J., Calvet, N., & Hartmann, L. 2001, ApJ, 550, 944
- Muzerolle et al. (1998b) Muzerolle, J., Hartmann, L., & Calvet, N. 1998b, AJ, 116, 455
- (279) NASA/ADS Developer API. , Online Guidelines:, https://ui.adsabs.harvard.edu/help/api/
- Natta et al. (2014) Natta, A., Testi, L., Alcalá, J. M., et al. 2014, A&A, 569, A5
- Natta et al. (2004) Natta, A., Testi, L., Muzerolle, J., et al. 2004, A&A, 424, 603
- Ochsenbein et al. (2000) Ochsenbein, F., Bauer, P., & Marcout, J. 2000, A&AS, 143, 23
- O’Dell et al. (2008) O’Dell, C. R., Muench, A., Smith, N., & Zapata, L. 2008, in Handbook of Star Forming Regions, Volume I, ed. B. Reipurth, Vol. 4, 544
- (284) OnlineOCR. , Online OCR Tool, https://www.onlineocr.net/
- Osuna et al. (2008) Osuna, P., Ortiz, I., Lusted, J., et al. 2008, IVOA Astronomical Data Query Language Version 2.00, IVOA Recommendation 30 October 2008
- Palla & Stahler (2001) Palla, F. & Stahler, S. W. 2001, ApJ, 553, 299
- pandas development team (2020) pandas development team, T. 2020, pandas-dev/pandas: Pandas
- Parihar et al. (2009) Parihar, P., Messina, S., Distefano, E., Shantikumar, N. S., & Medhi, B. J. 2009, MNRAS, 400, 603
- Peña Ramírez et al. (2012) Peña Ramírez, K., Béjar, V. J. S., Zapatero Osorio, M. R., Petr-Gotzens, M. G., & Martín, E. L. 2012, ApJ, 754, 30
- Pecaut & Mamajek (2013) Pecaut, M. J. & Mamajek, E. E. 2013, ApJS, 208, 9
- Pedregosa et al. (2011) Pedregosa, F., Varoquaux, G., Gramfort, A., et al. 2011, Journal of Machine Learning Research, 12, 2825
- Pettersson et al. (2014) Pettersson, B., Armond, T., & Reipurth, B. 2014, A&A, 570, A30
- Pickles & Depagne (2010) Pickles, A. & Depagne, É. 2010, PASP, 122, 1437
- Pilbratt et al. (2010) Pilbratt, G. L., Riedinger, J. R., Passvogel, T., et al. 2010, A&A, 518, L1
- Pillitteri et al. (2017) Pillitteri, I., Wolk, S. J., & Megeath, S. T. 2017, A&A, 608, L2
- Pillitteri et al. (2013) Pillitteri, I., Wolk, S. J., Megeath, S. T., et al. 2013, ApJ, 768, 99
- Pinzón et al. (2021) Pinzón, G., Hernández, J., Serna, J., et al. 2021, AJ, 162, 90
- Pittman et al. (2022) Pittman, C. V., Espaillat, C. C., Robinson, C. E., et al. 2022, AJ, 164, 201
- Popper (1993) Popper, D. M. 1993, PASP, 105, 721
- Popper & Plavec (1976) Popper, D. M. & Plavec, M. 1976, ApJ, 205, 462
- Povich et al. (2013) Povich, M. S., Kuhn, M. A., Getman, K. V., et al. 2013, ApJS, 209, 31
- Preibisch & Feigelson (2005) Preibisch, T. & Feigelson, E. D. 2005, ApJS, 160, 390
- Preibisch et al. (2005) Preibisch, T., Kim, Y.-C., Favata, F., et al. 2005, ApJS, 160, 401
- Proffitt et al. (2021) Proffitt, C. R., Roman-Duval, J., Taylor, J. M., et al. 2021, Research Notes of the American Astronomical Society, 5, 36
- Prša et al. (2022) Prša, A., Kochoska, A., Conroy, K. E., et al. 2022, ApJS, 258, 16
- (306) PS1 CasJobs. , PanSTARRS1 (PS1) Catalog Archive Server Jobs System (CasJobs) service., https://mastweb.stsci.edu/ps1casjobs/
- Qian et al. (2019) Qian, S.-B., Shi, X.-D., Zhu, L.-Y., et al. 2019, Research in Astronomy and Astrophysics, 19, 064
- Ramírez et al. (2004) Ramírez, S. V., Rebull, L., Stauffer, J., et al. 2004, AJ, 128, 787
- Randich et al. (2022) Randich, S., Gilmore, G., Magrini, L., et al. 2022, A&A, 666, A121
- Rebull (2001) Rebull, L. M. 2001, AJ, 121, 1676
- Rebull et al. (2000) Rebull, L. M., Hillenbrand, L. A., Strom, S. E., et al. 2000, AJ, 119, 3026
- Rebull et al. (2015) Rebull, L. M., Stauffer, J. R., Cody, A. M., et al. 2015, AJ, 150, 175
- Recio-Blanco et al. (2023) Recio-Blanco, A., de Laverny, P., Palicio, P. A., et al. 2023, A&A, 674, A29
- Reid et al. (1995) Reid, I. N., Hawley, S. L., & Gizis, J. E. 1995, AJ, 110, 1838
- Reipurth (2008a) Reipurth, B. 2008a, Handbook of Star Forming Regions, Volume I: The Northern Sky, Vol. 4
- Reipurth (2008b) Reipurth, B. 2008b, Handbook of Star Forming Regions, Volume II: The Southern Sky, Vol. 5
- Reipurth et al. (2007) Reipurth, B., Guimarães, M. M., Connelley, M. S., & Bally, J. 2007, AJ, 134, 2272
- Rhode et al. (2001) Rhode, K. L., Herbst, W., & Mathieu, R. D. 2001, AJ, 122, 3258
- Rice et al. (2015) Rice, T. S., Reipurth, B., Wolk, S. J., Vaz, L. P., & Cross, N. J. G. 2015, AJ, 150, 132
- Rieke et al. (2004) Rieke, G. H., Young, E. T., Engelbracht, C. W., et al. 2004, ApJS, 154, 25
- Riello et al. (2021) Riello, M., De Angeli, F., Evans, D. W., et al. 2021, A&A, 649, A3
- Rigliaco et al. (2011) Rigliaco, E., Natta, A., Randich, S., Testi, L., & Biazzo, K. 2011, A&A, 525, A47
- Rigliaco et al. (2012) Rigliaco, E., Natta, A., Testi, L., et al. 2012, A&A, 548, A56
- Rimoldini et al. (2023) Rimoldini, L., Holl, B., Gavras, P., et al. 2023, A&A, 674, A14
- Robberto et al. (2005) Robberto, M., Beckwith, S. V. W., Panagia, N., et al. 2005, AJ, 129, 1534
- Robberto et al. (2020) Robberto, M., Gennaro, M., Ubeira Gabellini, M. G., et al. 2020, ApJ, 896, 79
- Robitaille et al. (2006) Robitaille, T. P., Whitney, B. A., Indebetouw, R., Wood, K., & Denzmore, P. 2006, ApJS, 167, 256
- Rodrigo & Solano (2020) Rodrigo, C. & Solano, E. 2020, in XIV.0 Scientific Meeting (virtual) of the Spanish Astronomical Society, 182
- Rodrigo et al. (2012) Rodrigo, C., Solano, E., & Bayo, A. 2012, SVO Filter Profile Service Version 1.0, IVOA Working Draft 15 October 2012
- Rodríguez-Ledesma et al. (2009) Rodríguez-Ledesma, M. V., Mundt, R., & Eislöffel, J. 2009, A&A, 502, 883
- Roquette et al. (2020) Roquette, J., Alencar, S. H. P., Bouvier, J., Guarcello, M. G., & Reipurth, B. 2020, A&A, 640, A128
- Sacco et al. (2008) Sacco, G. G., Franciosini, E., Randich, S., & Pallavicini, R. 2008, A&A, 488, 167
- Samus’ et al. (2017) Samus’, N. N., Kazarovets, E. V., Durlevich, O. V., Kireeva, N. N., & Pastukhova, E. N. 2017, Astronomy Reports, 61, 80
- Sanchez et al. (2014) Sanchez, N., Inés Gómez de Castro, A., Lopez-Martinez, F., & López-Santiago, J. 2014, A&A, 572, A89
- Schiavon et al. (1997) Schiavon, R. P., Barbuy, B., Rossi, S. C. F., & Milone, A. 1997, ApJ, 479, 902
- Schlafly et al. (2019) Schlafly, E. F., Meisner, A. M., & Green, G. M. 2019, ApJS, 240, 30
- Schlieder et al. (2012) Schlieder, J. E., Lépine, S., Rice, E., et al. 2012, AJ, 143, 114
- Scholz & Eislöffel (2004) Scholz, A. & Eislöffel, J. 2004, A&A, 419, 249
- Scholz & Eislöffel (2005) Scholz, A. & Eislöffel, J. 2005, A&A, 429, 1007
- SEIP (2020) SEIP. 2020, Spitzer Enhanced Imaging Products
- Sergison et al. (2013) Sergison, D. J., Mayne, N. J., Naylor, T., Jeffries, R. D., & Bell, C. P. M. 2013, MNRAS, 434, 966
- Serna et al. (2021) Serna, J., Hernandez, J., Kounkel, M., et al. 2021, ApJ, 923, 177
- Sherry et al. (2004) Sherry, W. H., Walter, F. M., & Wolk, S. J. 2004, AJ, 128, 2316
- Sicilia-Aguilar et al. (2005) Sicilia-Aguilar, A., Hartmann, L. W., Szentgyorgyi, A. H., et al. 2005, AJ, 129, 363
- Silverberg et al. (2018) Silverberg, S. M., Kuchner, M. J., Wisniewski, J. P., et al. 2018, ApJ, 868, 43
- SIMBAD User’s Guide (v1.8) SIMBAD User’s Guide. v1.8, https://simbad.cds.unistra.fr/guide/otypes.htx
- Simon & Toraskar (2017) Simon, M. & Toraskar, J. 2017, ApJ, 841, 95
- Simon et al. (2004) Simon, T., Andrews, S. M., Rayner, J. T., & Drake, S. A. 2004, ApJ, 611, 940
- Skinner et al. (2003) Skinner, S., Gagné, M., & Belzer, E. 2003, ApJ, 598, 375
- Skrutskie et al. (2006) Skrutskie, M. F., Cutri, R. M., Stiening, R., et al. 2006, AJ, 131, 1163
- Slesnick et al. (2006) Slesnick, C. L., Carpenter, J. M., & Hillenbrand, L. A. 2006, AJ, 131, 3016
- Smith et al. (2023) Smith, G. D., Gillen, E., Hodgkin, S. T., et al. 2023, MNRAS, 523, 169
- Smith et al. (2005) Smith, N., Bally, J., Shuping, R. Y., Morris, M., & Kassis, M. 2005, AJ, 130, 1763
- Soderblom (2010) Soderblom, D. R. 2010, ARA&A, 48, 581
- Sousa et al. (2016) Sousa, A. P., Alencar, S. H. P., Bouvier, J., et al. 2016, A&A, 586, A47
- Spezzi et al. (2015) Spezzi, L., Petr-Gotzens, M. G., Alcalá, J. M., et al. 2015, A&A, 581, A140
- Spitzer Enhanced Imaging Products Explanatory Supplement (2013) Spitzer Enhanced Imaging Products Explanatory Supplement. 2013, Spitzer Enhanced Imaging Products Explanatory Supplement
- Stassun et al. (1999) Stassun, K. G., Mathieu, R. D., Mazeh, T., & Vrba, F. J. 1999, AJ, 117, 2941
- Stassun et al. (2006) Stassun, K. G., Mathieu, R. D., & Valenti, J. A. 2006, Nature, 440, 311
- Stassun et al. (2004) Stassun, K. G., Mathieu, R. D., Vaz, L. P. R., Stroud, N., & Vrba, F. J. 2004, ApJS, 151, 357
- Stassun & Torres (2021) Stassun, K. G. & Torres, G. 2021, ApJ, 907, L33
- Steinmetz et al. (2020) Steinmetz, M., Matijevič, G., Enke, H., et al. 2020, AJ, 160, 82
- Stelzer et al. (2007) Stelzer, B., Flaccomio, E., Briggs, K., et al. 2007, A&A, 468, 463
- Stempels et al. (2008) Stempels, H. C., Hebb, L., Stassun, K. G., et al. 2008, A&A, 481, 747
- Stetson (1987) Stetson, P. B. 1987, PASP, 99, 191
- Strampelli et al. (2020) Strampelli, G. M., Aguilar, J., Pueyo, L., et al. 2020, ApJ, 896, 81
- Struve (1945) Struve, O. 1945, ApJ, 102, 74
- Stutz et al. (2013) Stutz, A. M., Tobin, J. J., Stanke, T., et al. 2013, ApJ, 767, 36
- Suárez et al. (2019) Suárez, G., Downes, J. J., Román-Zúñiga, C., et al. 2019, MNRAS, 486, 1718
- Suárez et al. (2017) Suárez, G., Downes, J. J., Román-Zúñiga, C., et al. 2017, AJ, 154, 14
- Szegedi-Elek et al. (2013) Szegedi-Elek, E., Kun, M., Reipurth, B., et al. 2013, ApJS, 208, 28
- TableConvert (v 2.4.2) TableConvert. v 2.4.2, TableConvert
- (373) Tabula Tool. , Tool for table extraction from pdf files, https://tabula.technology/
- Taylor (2005) Taylor, M. B. 2005, in Astronomical Society of the Pacific Conference Series, Vol. 347, Astronomical Data Analysis Software and Systems XIV, ed. P. Shopbell, M. Britton, & R. Ebert, 29
- Thanathibodee et al. (2022) Thanathibodee, T., Calvet, N., Hernández, J., Maucó, K., & Briceño, C. 2022, AJ, 163, 74
- Theissen et al. (2022) Theissen, C. A., Konopacky, Q. M., Lu, J. R., et al. 2022, ApJ, 926, 141
- Theissen & West (2014) Theissen, C. A. & West, A. A. 2014, ApJ, 794, 146
- Tobin et al. (2009) Tobin, J. J., Hartmann, L., Furesz, G., Mateo, M., & Megeath, S. T. 2009, ApJ, 697, 1103
- Tobin et al. (2019) Tobin, J. J., Megeath, S. T., van’t Hoff, M., et al. 2019, ApJ, 886, 6
- Tokovinin et al. (2020) Tokovinin, A., Petr-Gotzens, M. G., & Briceño, C. 2020, AJ, 160, 268
- Traven et al. (2020) Traven, G., Feltzing, S., Merle, T., et al. 2020, A&A, 638, A145
- van Eyken et al. (2011) van Eyken, J. C., Ciardi, D. R., Rebull, L. M., et al. 2011, AJ, 142, 60
- Vasconcellos et al. (2011) Vasconcellos, E. C., de Carvalho, R. R., Gal, R. R., et al. 2011, AJ, 141, 189
- Venuti et al. (2015) Venuti, L., Bouvier, J., Irwin, J., et al. 2015, A&A, 581, A66
- Vioque et al. (2018) Vioque, M., Oudmaijer, R. D., Baines, D., Mendigutía, I., & Pérez-Martínez, R. 2018, A&A, 620, A128
- (386) VizieR. , The VizieR Catalogue Service, https://vizier.cds.unistra.fr/
- (387) VizieR Photometry viewer. , CDS VizieR-SED, http://vizier.cds.unistra.fr/vizier/sed/doc/, Accessed: 26 November 2024
- Walmsley et al. (2023) Walmsley, M., Géron, T., Kruk, S., et al. 2023, MNRAS, 526, 4768
- Walter et al. (1988) Walter, F. M., Brown, A., Mathieu, R. D., Myers, P. C., & Vrba, F. J. 1988, AJ, 96, 297
- Werner et al. (2004) Werner, M. W., Roellig, T. L., Low, F. J., et al. 2004, ApJS, 154, 1
- (391) WFCAM Science Archive. , , http://wsa.roe.ac.uk/dbaccess.html
- White & Basri (2003) White, R. J. & Basri, G. 2003, ApJ, 582, 1109
- Wilson et al. (2022) Wilson, T. J. G., Matt, S., Harries, T. J., & Herczeg, G. J. 2022, MNRAS, 514, 2162
- Windemuth et al. (2013) Windemuth, D., Herbst, W., Tingle, E., et al. 2013, ApJ, 768, 67
- Wright et al. (2010) Wright, E. L., Eisenhardt, P. R. M., Mainzer, A. K., et al. 2010, AJ, 140, 1868
- Wright et al. (2019) Wright, E. L., Eisenhardt, P. R. M., Mainzer, A. K., et al. 2019, AllWISE Source Catalog
- Yao et al. (2018) Yao, Y., Meyer, M. R., Covey, K. R., Tan, J. C., & Da Rio, N. 2018, ApJ, 869, 72
- Zari et al. (2018) Zari, E., Hashemi, H., Brown, A. G. A., Jardine, K., & de Zeeuw, P. T. 2018, A&A, 620, A172
- Zhao et al. (2012) Zhao, G., Zhao, Y.-H., Chu, Y.-Q., Jing, Y.-P., & Deng, L.-C. 2012, Research in Astronomy and Astrophysics, 12, 723
- Zheng et al. (2023) Zheng, Z., Cao, Z., Deng, H., et al. 2023, ApJS, 266, 18
Appendix A ALLWISE Cleaning
The WISE mission (Wright et al. 2010) was designed to target infrared-bright galaxies, brown dwarfs, and near-Earth asteroids. Hence, besides the convenience of its mid-infrared full-sky coverage, WISE observations of YSOs in embedded environments are often susceptible to spurious detections, especially towards the far-infrared. This issue has been previously investigated by a number of authors, and mitigation strategies based on AllWISE data release have been proposed to identify and remove affected sources (e.g. Koenig & Leisawitz 2014; Marton et al. 2016, 2019; Silverberg et al. 2018). In particular, Koenig & Leisawitz (2014, see their eq. 1 to 4) proposes that spurious detections can be separated by a series of cuts in terms of the signal-to-noise (w?snr) and profile fit reduced chi-squared (w?chi2) parameters in the AllWISE data release.
Here, we adopted a strategy similar to Marton et al. (2016, 2019), where we built a Random Forest (RF) classifier to evaluate the reliability of WISE W3 and W4 detections based on AllWISE quality parameters and magnitudes. Nevertheless, we introduced two main modifications to Marton et al. (2019) approach. First, Marton et al. (2019) evaluated the reliability of W3 and W4 photometry simultaneously. Contrarily, we expect our sample to include sources that may only have one of these two bands above WISE’s detection limits - thus we have trained one RF classifier for each photometric band. Second, Marton et al. (2019) classifier was trained with a sample of labelled AllWISE cutouts, where a researcher visually inspected the presence or lack of a source in image cutouts at a given band. Our initial tests with this approach revealed that such labels were highly subject to the researcher carrying out the labelling. To remove this subjectivity, we used instead Astropy’s DAOStarFinder (Stetson 1987; Bradley et al. 2024) to search the cutout images for local density maxima and label them in case a source detection was successful. We have built a training sample based on 5,000 AllWISE cutout images around known YSOs in Orion. These cutouts have been obtained for each of the W3 and W4 bands through the IRSA WISE Image Service and covered a box of 10” around the positions of the sources. Our DAOStarFinder approach allowed us to label 2,356 and 1,970 sources as true detections in the W3 and W4 bands respectively.
| label | precision | recall | f1-score |
| W3 | |||
| real | 0.98 | 0.98 | 0.98 |
| false | 0.97 | 0.97 | 0.97 |
| W4 | |||
| real | 0.90 | 0.91 | 0.90 |
| false | 0.90 | 0.89 | 0.90 |
Our RF classifiers999We experimented with various machine learning classification methods, including also Logistic Regression, k-Neighbours, Naive Bayes, Decision Tree, Support Vector Machine, and Neural Networks, but found that the RF classifier yields the highest classification accuracy during the testing phase. were trained on samples of 3,800 labelled sources (with 50-50% labelled as true and false detections), using 21 features per band: AllWISE profile-fit magnitudes and their uncertainties (w?mpro, w?sigmpro), signal-to-noise (w?snr) and profile fit reduced chi-squared (w?chi2), the magnitudes and uncertainties at 8 different apertures (w?mag_# and w?sigm_#, where # refer to the 8 available apertures within AllWISE), and the ratios between magnitudes at consecutive apertures. Classification reports for both bands are given in Table 2 and Confusion matrices are shown in Fig. 14.
Next, we have employed the trained RF classifiers to estimate reliability probabilities for all available W3 and W4 AllWISE data in the OSFC field. This allowed us to flag 1,760,122 and 660,689 AllWISE observations in the W3 and W4 bands respectively. The distribution of reliability probabilities for this sample is shown in Fig. 15. For our purpose in this paper, we have removed all W3 and W4 data that had probability inferior to 60%. Given the sources in our compilation with AllWISE counterparts, 17,453 (W3) and 6,581 (W4) sources had reliability probability derived, out of which 10,688 (W3) and 4,529 (W4) were under the threshold. Additionally, 404 (W3) and 230 (W4) sources had AllWISE counterparts in these bands but did not have all features used by the classifier available and we also eliminated these from the compilation.
Appendix B reliability flags
To quantify the effect of the reduced availability of mid/far infrared data on our -indices (Sect. 3.3.1), we carried out an undersampling exercise. Recalling our requirement of two data points at least 2 away in wavelength for derivation, we selected a testing sample of 16 282 sources that followed this requirement for wavelengths m, and had at least one additional data point in the range m. For this sample, we derived an additional -index using only the data between m. Fig. 16 shows a comparison of this index to our main . For the specific discussion in this Appendix, we limit our classes to three: discless, () thin disc () and thick disc (). We find that of all sources had the same class with either definition (white and gray-shaded areas in Fig. 16); sources were discless sources miss-detected as thin disc (yellow); were thick discs miss-detected as thin disc (green); were thin discs miss-detected as thick disc (red); of sources were thin discs miss-detected as discless (purple); were thick discs miss-detected as discless (blue); and sources were discless miss-detected as thick discs (cyan). With 33% of our YSO candidates having no SED datapoints for m, we expected such bias to affect of our -indices. We note that this estimation does not take into account domain-motivated choices of the wavelength range used for , and we point the interested reader to Großschedl et al. (2019) for an in-depth discussion of the effect of definitions on disc classification. We also note that the numbers discussed so far refer to the total number of sources analysed. In the bottom plot of Fig. 16, we detailed miss-classification rates normalised by the number of sources in each class derived with . Although only 71.2 and 66.9% of the original thick discs and diskless sources are recovered as such, these classes have relatively large purity, with 93.7% of thick disks and 97.8% of discless sources correctly detected as such with the index. In contrast, while 88.6% of thin discs are recovered as such, this class has a much smaller purity of 74.8%.
Appendix C Summary of studies included in the historical compilation
Tab. LABEL:tab:papers_with_data presents a summary of peer-reviewed scientific publications with data collated into the NEMESIS catalogue of YSOs for the Orion Star Formation Complex.
1]
| Reference | Data type | CDS | pub | man | TbC | Tba |
| Struve (1945) | bin | x | ||||
| Popper & Plavec (1976) | bin | x | ||||
| Hesser et al. (1976) | bin | x | ||||
| Marschall & Mathieu (1988) | bin | x | ||||
| Bossi et al. (1989) | bin | x | ||||
| Mathieu et al. (1991) | bin | x | ||||
| Abt et al. (1991) | bin | x | ||||
| Popper (1993) | bin | x | ||||
| Gagne & Caillault (1994) | X-ray | x | ||||
| Kozlova et al. (1995) | SpT, EWacc, H | x | x | |||
| Gies et al. (1996) | bin | x | ||||
| Hillenbrand (1997) | , SED | x | ||||
| Bolton et al. (1998) | bin | x | ||||
| Hillenbrand et al. (1998) | EWacc disc, SED, | x | ||||
| Stassun et al. (1999) | Li, RV, H, SED | x | ||||
| Lloyd & Stickland (1999) | bin | x | ||||
| Alcalá et al. (2000) | X-ray, Li, rotation, RV, SpT, H, bin | x | x | x | ||
| Covino et al. (2000) | bin | x | ||||
| Herbst et al. (2000) | rotation, EWacc | x | x | |||
| Rebull et al. (2000) | SpT, SED | x | ||||
| Covino et al. (2001) | bin | x | ||||
| Rebull (2001) | rotation, SpT, SED | x | ||||
| Dolan & Mathieu (2001) | Li, RV, H, bin, SED | x | ||||
| Carpenter et al. (2001) | rotation, SED | x | ||||
| Rhode et al. (2001) | rotation, bin | x | ||||
| Palla & Stahler (2001) | bin | x | ||||
| Béjar et al. (2001) | SpT, SED | x | ||||
| Kharchenko (2001) | SpT, bin | x | ||||
| Herbst et al. (2002) | rotation | x | ||||
| Manset & Bastien (2002) | bin | x | ||||
| Feigelson et al. (2002) | X-ray | x | ||||
| Maheswar et al. (2003) | Li, SpT, H, SED | x | x | x | ||
| Flaccomio et al. (2003) | X-ray | x | ||||
| Skinner et al. (2003) | X-ray | x | ||||
| Alcalá et al. (2004) | Li, RV, , SpT, H, bin | x | x | |||
| Scholz & Eislöffel (2004) | rotation , SpT, logg, EWacc, H, SED | x | x | x | x | |
| Ramírez et al. (2004) | X-ray | x | ||||
| Sherry et al. (2004) | SED | x | ||||
| Barrado y Navascués et al. (2004) | SpT, H, SED | x | ||||
| Simon et al. (2004) | X-ray | x | x | |||
| Ali & Noriega-Crespo (2004) | disc, SED | x | x | |||
| Stassun et al. (2004) | Li, bin | x | ||||
| Baines et al. (2004) | bin | x | ||||
| Scholz & Eislöffel (2005) | rotation, SED | x | x | x | ||
| Sicilia-Aguilar et al. (2005) | Li, rotation, RV, SpT, H, bin | x | ||||
| Briceño et al. (2005) | Li, , SpT, H | x | ||||
| Robberto et al. (2005) | bin, SED | x | ||||
| Smith et al. (2005) | disc, SED | x | ||||
| Getman et al. (2005b) | X-ray | x | ||||
| Getman et al. (2005a) | Source list | |||||
| Kenyon et al. (2005) | Li, RV, EWgravity, bin | x | x | |||
| Burningham et al. (2005) | RV, EWgravity, SED | x | ||||
| Franciosini et al. (2006) | X-ray, SpT | x | ||||
| Köhler et al. (2006) | bin | x | ||||
| Herbig & Griffin (2006) | bin | x | ||||
| Stassun et al. (2006) | bin | x | ||||
| Caballero (2007) | RV, SpT, SED | x | ||||
| Barrado Y Navascués et al. (2007) | bin | x | ||||
| Reipurth et al. (2007) | SpT, bin | x | ||||
| Lee & Chen (2007) | SpT, H | x | x | |||
| Briceño et al. (2007) | Li, RV, H | x | ||||
| Hernández et al. (2007b) | disc, SED | x | ||||
| Barrado y Navascués et al. (2007) | disc, SED | x | ||||
| Hernández et al. (2007a) | disc, SED | x | x | |||
| Irwin et al. (2007) | bin | x | ||||
| Caballero (2008) | SED | x | ||||
| Stempels et al. (2008) | bin | x | ||||
| Caballero & Solano (2008) | SpT, bin, SED | x | ||||
| Sacco et al. (2008) | Li, rotation, RV, SpT, EWacc, H, bin | x | ||||
| López-Santiago & Caballero (2008) | X-ray | x | x | |||
| Bayo et al. (2008) | x | |||||
| Cargile et al. (2008) | bin | x | ||||
| Fűrész et al. (2008) | RV, H | x | ||||
| Luhman et al. (2008b) | SED | x | ||||
| Maxted et al. (2008) | RV, EWgravity, bin, SED, | x | ||||
| Rodríguez-Ledesma et al. (2009) | rotation, SED, | x | ||||
| Fang et al. (2009) | Li, SpT, EWacc, H, disc, SED, | x | x | |||
| Biazzo et al. (2009) | rotation, RV, SpT, H, | x | x | x | ||
| Frasca et al. (2009) | rotation, , SpT, SED, | x | ||||
| Mace et al. (2009) | bin, | x | ||||
| Baldovin-Saavedra et al. (2009) | bin, | x | ||||
| Mohanty et al. (2009) | bin, | x | ||||
| Tobin et al. (2009) | RV, bin, SED, | x | ||||
| Mookerjea & Sandell (2009) | disc, SED, | x | ||||
| Hernández et al. (2009) | disc, SED, | x | ||||
| Da Rio et al. (2009) | SpT, H, SED, | x | ||||
| Alecian et al. (2009) | bin, | x | ||||
| Parihar et al. (2009) | rotation, H, SED, | x | ||||
| Connelley & Greene (2010) | SpT, EWacc, veiling, bin, disc, SED, | x | ||||
| Grinin et al. (2010) | bin, | x | ||||
| Da Rio et al. (2010) | , SpT, | x | ||||
| Cody & Hillenbrand (2010) | rotation, SED, | x | ||||
| Leone et al. (2010) | bin, | x | ||||
| Rigliaco et al. (2011) | , H, disc, SED, | x | x | |||
| Barrado et al. (2011) | X-ray, disc, SED, | x | ||||
| Franciosini & Sacco (2011) | X-ray, SpT, disc, | x | ||||
| Bayo et al. (2011) | Li, , SpT, EWgravity, | x | ||||
| Ingleby et al. (2011) | X-ray, SpT, H, | x | x | |||
| van Eyken et al. (2011) | bin, | x | ||||
| Morales-Calderón et al. (2011) | rotation, | x | ||||
| Béjar et al. (2011) | disc, SED, | x | ||||
| Daemgen et al. (2012) | bin, | x | ||||
| Caballero et al. (2012) | Li, SpT, EWacc, H, disc, | x | x | |||
| Bayo et al. (2012) | rotation, SpT, EWacc, H, disc, | x | x | |||
| Rigliaco et al. (2012) | SpT, EWacc, H, | x | ||||
| Kryukova et al. (2012) | disc, SED, | x | ||||
| Megeath et al. (2012) | disc, SED, | x | ||||
| Gómez Maqueo Chew et al. (2012) | bin, | x | ||||
| Da Rio et al. (2012) | , | x | ||||
| Hsu et al. (2012) | Li, SpT, EWacc, H, SED, | x | ||||
| Morales-Calderón et al. (2012) | rotation, SpT, bin, | x | ||||
| Peña Ramírez et al. (2012) | SED, | x | ||||
| Manara et al. (2012) | H, disc, | x | ||||
| Correia et al. (2013) | bin, | x | ||||
| Hillenbrand et al. (2013) | SpT, bin, | x | ||||
| Hsu et al. (2013) | Li, RV, SpT, disc, SED, | x | ||||
| Stutz et al. (2013) | , bin, disc, SED, | x | ||||
| Windemuth et al. (2013) | bin, | x | ||||
| Pillitteri et al. (2013) | X-ray, | x | ||||
| Kim et al. (2013) | X-ray, , SpT, bin, disc, | x | ||||
| Fang et al. (2013) | Li, SpT, EWacc, H, disc, SED, | x | ||||
| Szegedi-Elek et al. (2013) | H, SED, | x | ||||
| Kuhn et al. (2013) | SED, | x | ||||
| Povich et al. (2013) | Source list, | |||||
| Broos et al. (2013) | X-ray, disc, | x | ||||
| López-García et al. (2013) | X-ray, | x | ||||
| Sergison et al. (2013) | Li, RV, veiling, H, | x | ||||
| Bouy et al. (2014) | SED, | x | ||||
| Pettersson et al. (2014) | H, bin, SED, | x | ||||
| Fang et al. (2014) | bin, | x | ||||
| Sanchez et al. (2014) | , SED, | x | ||||
| Ingraham et al. (2014) | , SpT, | x | ||||
| Kuhn et al. (2014) | SpT, bin, | x | ||||
| Hernández et al. (2014) | Li, RV, SpT, EWgravity, H, bin, disc, SED, | x | x | x | ||
| Theissen & West (2014) | RV, , SpT, EWgravity, EWacc, H, | x | ||||
| Downes et al. (2014) | SpT, H, | x | ||||
| Spezzi et al. (2015) | disc, SED, | x | ||||
| Koenig et al. (2015) | SpT, EWacc, H, disc, SED, | x | ||||
| Rice et al. (2015) | rotation, bin, SED, | x | ||||
| Downes et al. (2015) | , SpT, H, disc, SED, | x | ||||
| Meingast et al. (2016) | SED, | x | ||||
| Hasenberger et al. (2016) | SpT, disc, | x | ||||
| Megeath et al. (2016) | disc, SED, | x | ||||
| Karim et al. (2016) | rotation, | x | ||||
| Da Rio et al. (2016) | rotation, , logg, SED, | x | ||||
| Kounkel et al. (2016a) | Li, RV, , bin, | x | ||||
| Kounkel et al. (2016b) | bin, disc, | x | ||||
| Lewis & Lada (2016) | bin, disc, | x | ||||
| Furlan et al. (2016) | , disc, SED, | x | ||||
| Kim et al. (2016) | Li, , SpT, H, bin, disc, | x | ||||
| Messina et al. (2016) | bin, | x | ||||
| Pillitteri et al. (2017) | X-ray, | x | x | |||
| Fang et al. (2017) | Li, SpT, H, disc, | x | ||||
| Suárez et al. (2017) | Li, , SpT, H, disc, | x | ||||
| Kounkel et al. (2017a) | Li, SpT, H, | x | ||||
| Kounkel et al. (2017b) | bin, | x | ||||
| Simon & Toraskar (2017) | bin, | x | ||||
| Kounkel et al. (2017c) | Li, RV, , H, disc, | x | ||||
| Jaehnig et al. (2017) | bin, | x | ||||
| Getman et al. (2017) | X-ray, SED, | x | ||||
| Messina et al. (2017) | SED, | x | ||||
| Fernandez et al. (2017) | bin, | x | ||||
| GRAVITY Collaboration et al. (2018) | bin, | x | ||||
| Vioque et al. (2018) | , H, bin, disc, SED, | x | ||||
| Zari et al. (2018) | RV, | x | ||||
| Kounkel et al. (2018) | RV, , logg, | x | ||||
| Caballero et al. (2018) | bin, | x | ||||
| Grant et al. (2018) | SpT, disc, SED, | x | ||||
| Yao et al. (2018) | , logg, EWacc, disc, SED, | x | ||||
| Cottle et al. (2018) | disc, SED, | x | ||||
| Davies et al. (2018) | bin, | x | ||||
| Getman et al. (2018) | bin, | x | ||||
| Getman et al. (2018) | disc, | |||||
| Jayasinghe et al. (2018) | rotation, | x | ||||
| Duchêne et al. (2018) | bin, | x | ||||
| Großschedl et al. (2019) | disc, SED, | x | ||||
| Jerabkova et al. (2019) | bin, | x | ||||
| Caballero et al. (2019) | Li, SpT, EWacc, H, | x | x | x | ||
| Briceño et al. (2019) | Li, , SpT, EWgravity, H, SED, | x | ||||
| Kim et al. (2019) | bin, SED, | x | ||||
| Kounkel et al. (2019) | RV, , logg, veiling, bin, disc, | x | ||||
| Mairs et al. (2019) | bin, | x | ||||
| McBride & Kounkel (2019) | RV, | x | ||||
| Tobin et al. (2019) | bin, | x | ||||
| De Furio et al. (2019) | bin, | x | ||||
| Suárez et al. (2019) | SED, | x | ||||
| Bouma et al. (2020) | bin, | x | ||||
| Tokovinin et al. (2020) | bin, | x | ||||
| Kounkel et al. (2020) | Source list, | |||||
| Manzo-Martínez et al. (2020) | SpT, H, | x | ||||
| Robberto et al. (2020) | , | x | ||||
| Strampelli et al. (2020) | bin, | x | ||||
| Fischer et al. (2020) | SED, | x | ||||
| Jackson et al. (2020) | RV, , logg, SED, | x | ||||
| Lavail et al. (2020) | bin, | x | ||||
| Frasca et al. (2021) | Source list, | |||||
| Pinzón et al. (2021) | rotation, RV, , SpT, bin, disc, | x | x | |||
| Fang et al. (2021) | SpT, veiling, bin, | x | ||||
| Habel et al. (2021) | disc, | x | ||||
| Laos et al. (2021) | EWacc, | x | x | x | ||
| Serna et al. (2021) | rotation, , bin, SED, | x | ||||
| Arun et al. (2021) | bin, | x | ||||
| Kos et al. (2021) | Li, rotation, RV, , logg, | |||||
| Proffitt et al. (2021) | bin, | x | ||||
| Franciosini et al. (2022) | RV, , | x | ||||
| Flaischlen et al. (2022) | , | x | ||||
| Thanathibodee et al. (2022) | , SpT, EWacc, H, bin, | x | ||||
| Kounkel et al. (2022) | rotation, | x | ||||
| Pittman et al. (2022) | rotation, | x | ||||
| Cao et al. (2022) | , disc, | x | ||||
| De Furio et al. (2022b) | bin, | x | ||||
| Theissen et al. (2022) | rotation, RV, , veiling, bin, | x | ||||
| De Furio et al. (2022a) | bin, | x | ||||
| Hernández et al. (2023) | Li, SpT, H, | x | ||||
| Federman et al. (2023) | , disc, | x | ||||
| Damian et al. (2023) | , SpT, SED, | x | ||||
| Smith et al. (2023) | rotation, , disc, | x |
Appendix D Database Summary
The database curated as part of the NEMESIS Catalogue of YSOs for the OSFC can be accessed via SQL (at https://www.astro.unige.ch/nemesis/) and through CDS. The 19 thematic tables with data compiled in the present study are organised in 12 directories. A full description of directories, sub-tables, column names, units and data availability is also available as part of the online material (nemesis_osfc_description.csv). This appendix further summarises key data products, the number of individual measurements collected for each observable, and the number of sources with at least one measurement. We stress that the data is organised following a relational database paradigm. Each source is associated with a unique NEMESIS_ID key. This unique source identifier should be used to group multiple measurements (when available) for a given source and match different thematic tables.
For each quantity described in Tabs. to , when available, the catalogue includes the following accompanying fields:
#: reported quantity, for example Teff for or EW_Ha for equivalent width of the H line;
#_ref: NASA/ADS bibcode for the original paper linked to the reported quantity #;
e_# uncertainty;
e_up_#/e_low_# upper and lower uncertainty values;
f_# flag indicating that a limit value is reported;
#_comment: String reporting author’s comments from the original paper.
Tab. ir/ : Quantities related to IR YSO classification schemes discussed Sect. 3.3;
standard_classification_nemesis.csv YSO classification derived in this study.
disk_literature.csv: IR-based YSO classification collected from the literature.
Tab. : accretion/ : Quantities related to accretion or material inflow/outflow (Sect. 3.6);
Tab. : quantities related to the X-ray emission of YSOs (Sect. 3.9);
Tab. multiplicity/multiplicity_label.csv: Multiplicity labels (Sect. 4).
| key | description | units | count | # sources |
| Identification (main/id.csv) | ||||
[…]
| Internal_ID | NEMESIS Catalogue Internal ID | - | 27 879 | 27 879 |
| Simbad counterparts (main/simbad.csv) | ||||
| main_id | Main identifier in Simbad | - | 33 846 | 25 902 |
| otype | Main Object Type in Simbad | - | 33 690 | 25 901 |
| otypes | Full list of Object types in Simbad | - | 33 615 | 25 902 |
| ra_simbad, dec_simbad | RA, Dec coordinates in Simbad | degree | 33 846 | 25 902 |
[…]
| counterpart_groupID | identifier for grouping multiple possible counterparts | - | 33 846 | 25 902 |
| sep2simbad | Separation to Simbad counterpart | arcsec | 12 835 | 4 891 |
| Contamination (main/contamination.csv)) | ||||
| massive_star | Flag indicative of likely massive star (Sect. 5.2.1) | - | 80 | 80 |
| ms_contaminant | Likely MS contaminant(Sect. 5.2.4) | - | 2 845 | 2 845 |
| extraglactic_prob | Probability of source being extragalactic (App. E) | - | 15 855 | 15 855 |
| extraglactic_pred | Extragalactic predicted label (App. E) | - | 15 855 | 15 855 |
| extraglactic_label | Labelled extragalactic contaminant (Sect. 5.2.3 and App. E.2.1) | - | 125 | 125 |
| giant_prob | Probability of source being a Giant (App. E) | - | 20 546 | 20 546 |
| giant_prob | Probability of source being a Giant (App. E) | - | 20 546 | 20 546 |
| giant_label | Labelled Giant contaminant (App. E.2.1 and Sect. 5.2.2) | - | 490 | 490 |
[…]
| key | description | units | count | # sources |
| Temperature (hrd/temperature.csv) | ||||
| Teff | Effective temperature | K | 73 808 | 17 589 |
[…]
| Tbol | Bolomnetric temperature | K | 700 | 381 |
| Spectral Type (hrd/spt.csv) | ||||
[…]
| phot_SpT | True if Spectral typing was based on narrow-band photometry. | - | 263 | 263 |
| field | description | units |
[…]
| nu_Fnu | observed flux density | |
| nu_Fnu_error | flux density uncertainty | |
| wavelength | effective wavelength | Å |
| flux | observed flux | Jy |
| freq | frequency | Hz |
| bib | NASA/ADS bibcode for the original paper linked to the reported quantity | - |
| filter | SVO filter name | - |
| SED summary table (sed/sed_summary.csv) | ||
| uv | Number of SED datapoints in the range Å | - |
| optical_1 | Number of SED datapoints in the range Å | - |
| optical_2 | Number of SED datapoints in the range Å | - |
| optical_3 | Number of SED datapoints in the range Å | - |
| nir_1 | Number of SED datapoints in the range Å | - |
| nir_2 | Number of SED datapoints in the range Å | - |
| nir_3 | Number of SED datapoints in the range Å | - |
| mir_1 | Number of SED datapoints in the range Å | - |
| mir_2 | Number of SED datapoints in the range Å | - |
| mir_3a | Number of SED datapoints in the range Å | - |
| mir_3b | Number of SED datapoints in the range Å | - |
| mir_4 | Number of SED datapoints in the range Å | - |
| mir_5 | Number of SED datapoints in the range Å | - |
| fir_1 | Number of SED datapoints in the range Å | - |
| fir_2 | Number of SED datapoints in the range Å | - |
| fir_3 | Number of SED datapoints in the range Å | - |
| key | description | units | count | # sources |
| Standard Classification Scheme (ir/stnadard_classification_nemesis.csv) | ||||
| alpha_2_24 | estimated in Sect. 3.3.1. | - | 25 799 | 25 799 |
[…]
| lambda_min_2_24 | Minimum wavelength available for estimation | m | 25 799 | 25 799 |
| lambda_max_2_24 | Maximum wavelength available for estimation | m | 25 799 | 25 799 |
| Literature values (disk_literature.csv) | ||||
| alpha | compiled from the literature | - | 48 657 | 5 379 |
| class | IR class compiled from the literature | - | 16 716 | 8 097 |
| IR_range | wavelength range used for classification | - | 62 986 | 9 438 |
[…]
| disc | disc description keyword | - | 8 054 | 4 096 |
[…]
| key | description | units | count | # sources |
| Lithium (lithium/li.csv) | ||||
| EW_Li | Equivalent Width at 6708 Å | Å | 8 906 | 6 337 |
| A_Li | Lithium abundance | dex | 2 324 | 1 997 |
| key | description | units | count | # sources |
| (gravity/logg_DB.csv) | ||||
| logg | derived from spectra fit | dex | 57 649 | 12 313 |
| line measurements (gravity/ew_gravity.csv) | ||||
| Na I doublet in absorption | ||||
| NaI_8200 | Equivalent width at 8200 Å(unresolved doublet) | Å | 1 451 | 1 362 |
| NaI_8195 | Equivalent width at 8195 Å | Å | 102 | 102 |
| K I doublet in absorption | ||||
| EW_KI_7665 | Equivalent width at 7665 Å | Å | 52 | 52 |
| EW_KI_7699 | Equivalent width at 7699 Å | Å | 39 | 39 |
| EW_KI_7665_7699 | Equivalent width at 7665-7699 Å | Å | 6 | 6 |
| CaH3 | ||||
| EW_CaH3_6965 | Equivalent width at 6965 Å | Å | 6 | 6 |
| TiO | ||||
| EW_TiO_8442 | Equivalent width at 8442 Å | Å | 218 | 218 |
| TiO5 | ||||
| EW_TiO_7130 | Equivalent width at 7130 Å | Å | 6 | 6 |
| key | description | units | count | # sources |
| H (accretion/Ha_DB.csv) | ||||
| EW_Ha | EW of at Å | Å | 32 760 | 20 667 |
| FW10%(H) | Full-width at 10 of the H emission profile peak | km/s | 1 181 | 945 |
| EW_Ha_PHOT | EW of from photometry | Å | 1 514 | 1 239 |
[…]
| other lines(accretion/EW_accretion_other_DB.csv) | ||||
| Other Hydrogen Lines | ||||
| EW_Hb_4861 | EW of H at Å | Å | 2357 | 2319 |
| EW_Hg_4340 | EW of H at Å | Å | 57 | 48 |
| EW_H11_3771 | EW of H11 at Å | Å | 8 | 8 |
| EW_Pab_12822 | EW of Paschen at Å | Å | 15 | 15 |
| EW_Pag | EW of Pa ( Å) | Å | 7 | 7 |
| EW_Brg_21661 | EW of Brackett at Å | Å | 29 | 29 |
| EW_Br11_16810 | EW of Brackett 11 at Å | Å | 158 | 158 |
| Helium Lines | ||||
| EW_HeI_6678 | EW of He I line at 6678 Å | Å | 249 | 231 |
| EW_HeI_5876 | EW of He I line at 5876 Å | Å | 63 | 57 |
| EW_HeI_10830 | EW of He I line at 10830 Å | Å | 58 | 54 |
| Calcium Lines | ||||
| EW_CaII_8498 | EW of Ca II triplet at 8498 Å | Å | 130 | 122 |
| EW_CaII_8542 | EW of Ca II triplet at 8542 Å | Å | 1142 | 1134 |
| EW_CaII_8662 | EW of Ca II triplet at 8662 Å) | Å | 135 | 129 |
| EW_CaII_H_3934 | EW of Ca II HK doublet at 3964 Å | Å | 49 | 41 |
| EW_CaII_K_3969 | EW of Ca II HK doublet at 3969 Å | Å | 53 | 45 |
| Forbidden Lines | ||||
| EW__OI__5577 | EW of [O I] at 5577 Å | Å | 72 | 69 |
| EW__OI__6300 | EW of [O I] at 6300 Å | Å | 189 | 173 |
| EW__OI__6364 | EW of [O I] at 6364 Å | Å | 75 | 69 |
| EW__SII__6716 | EW of [S II] at 6717Å | Å | 162 | 157 |
| EW__SII__6731 | EW of [S II] at 6731Å | Å | 169 | 160 |
| EW__FeII__16440 | EW of [Fe II] at 16440Å | Å | 6 | 6 |
| EW__NII__6548 | EW of [N II] at 6548Å | Å | 18 | 18 |
| EW__NII__6581 | EW of [N II] at 6581Å | Å | 16 | 16 |
| EW__NII__6583 | EW of [N II] at 6583Å | Å | 128 | 128 |
| Other emission lines | ||||
| EW_OI_7773 | EW of the OI at 7773 Å | Å | 49 | 49 |
| EW_OI_8446 | EW of the OI line at 8446 Å | Å | 54 | 54 |
| EW_NaI_5893 | EW of the NaI doublet at 5893 Å | Å | 8 | 8 |
| EW_H2_1_0_S(1) | EW of the H2 1-0 S(1) emission at 2.1218m | Å | 13 | 13 |
| EW_CO_2_0 | EW of the CO 2-0 1st overtone at m | Å | 24 | 24 |
| Veiling (accretion/veiling_DB.csv) | ||||
| veiling_r_opt | optical-veiling at 6390-6710Å | - | 779 | 779 |
| veiling_r_h | ir-veiling at the H-band (m) | - | 32 132 | 9 031 |
| veiling_r_k | ir-veiling at the K-band range (m) | - | 3 | 3 |
| veiling_r_r7465 | veiling around 7465Å | - | 1 444 | 361 |
| key | description | units | count | # sources |
| RV | RV | km/s | 60 294 | 13 721 |
| key | description | units | count | # sources |
| Variability (variability/amplitude.csv) | ||||
| AG_proxy | Gaia DR3 -proxy variability index | - | 24 632 | 24 632 |
| key | description | units | count | # sources |
| Rotation (rotation/rotation.csv) | ||||
| Per | Variability periods | day | 21 932 | 10 147 |
| Per_band | Photometric band used for period measurement | - | - | - |
| vsini | km/s | 20 820 | 10 651 | |
| key | description | unit | count | # sources |
| xray/Xray_DB.csv | ||||
| Fx | X-ray flux | erg/cm2/s | 2 371 | 2 333 |
| logLx | Total X-ray luminosity | erg/s | 6 121 | 3 051 |
| #_band | Energy band observed. # paired with Fx or logLx. | keV | ||
| logNH | Hydrogen column density | 4 324 | 1 952 | |
| HR# | Hardness ratio #, # in (1, 2, or 3) | - | 1 364 | 1 364 |
[…]
| key | description | unit | count | # sources |
| multiplicity/multiplicity_label.csv | ||||
[…]
| bin_ext_type | Multiplicity type in source paper | - | 18 659 | 12 102 |
Appendix E Extragalactic and Giants contamination assessment
To investigate extragalactic and giant contamination in our catalogue, we used Scikit-learn (Pedregosa et al. 2011) to design two binary Random Forest classifiers (Breiman 2001) trained to disaggregate these types of sources based on a set of attributes and a labelled sample. Similar classifiers specialised in the separation of stellar and extragalactic sources have been extensively implemented in the literature (e.g., Vasconcellos et al. 2011; Logan & Fotopoulou 2020; Cook et al. 2024; Lourens et al. 2024). However, these previous studies do not explicitly include YSOs in their training sample. Hence, here, we build our own classifier tailored for the OSFC field.
E.1 Data
We built a reference photometric catalogue for the OSFC field based on large photometric surveys summarised in Sect. 2.2, with a base list of 16 177 367 sources with data from Gaia DR3, PanSTARRS1-DR2 and CatWISE. Near-infrared data were excluded because there were no surveys covering the entire OSFC field with comparable depth to the other surveys. We refer to the RF classifier trained to identify extragalactic sources against stellar sources as “Extragalactic classifier”, and the one trained to identify giant stars against stellar and extragalactic sources, as “Giant classifier”.
E.2 Labelled dataset
E.2.1 Extragalactic sources
We collated a sample of previously identified extragalactic sources, including a mixture of resolved and unresolved Galaxies, AGNs and QSOs. This accounts for 32 850 sources in the OSFC field. To maximise purity, we kept only extragalactic sources with a match within to the photometric reference catalogue. The list of extragalactic sources was composed of sources from:
HyperLEDA database (Makarov et al. 2014): 2 638 sources;
Million Quasars catalogue v. 8 (Flesch 2023): 4 641 sources;
Galaxy Zoo (Walmsley et al. 2023): 6 124 sources;
MANGROVE Catalogue (Biteau 2021): 1 396 sources;
Gaia DR3 high-purity Galaxy sample: 7 718 sources;
Gaia DR3 high-purity QSO sample: 15 629 sources.
Extragalactic sources from Gaia DR3 were identified by Gaia’s ML classification and redshift estimation (Delchambre et al. 2023), brightness profile analysis (Ducourant et al. 2023), or variability behaviour (Carnerero et al. 2023; Rimoldini et al. 2023). High-purity samples were devised based on selection procedures outlined by Gaia Collaboration et al. (2023b).
E.2.2 Stellar sources
The following samples of stellar sources were included in our labelled sample:
Certainly young: 3 698 sources from the NEMESIS YSO catalogue which have spectroscopic confirmation of youth based on significant Lithium absorption in their spectra (Sect. 3.4);
At Orion’s distance: 59 146 Gaia DR3 sources with parallax consistent with a location around the location of the OSFC (300-500 pc) and parallax uncertainty better than 0.05 mas;
Giants: 13 336 sources with in either RAVE, LAMOST LRS or MRS (Sect. 2.3), and precision better than 0.25 dex;
Other Stellar Sources: 72 336 sources with in either RAVE, LAMOST LRS or MRS (Sect. 2.3), and precision better than 0.25 dex.
E.2.3 Final sample
The labelled data were further processed to remove any overlapping labels. For example, sources labelled both Certainly Young and Giant were removed. Further pre-cleaning of the data included removal of sources with very large uncertainties.
Extragalactic classifier sample:
All sources listed in Sect. E.2.1 were initially labelled as “extragalactic” (positive label), and all sources in Sect. E.2.2 were labelled as “stellar” (negative label). Validation of labelled samples was carried out by comparing the locus occupied by the different samples in colour-colour diagrams. During this procedure, we verified a significant tail of “Gaia DR3 high-purity QSOs” sources towards the location of stellar sources, which did not exist in the distribution of other samples with large numbers of QSOs (e.g., Million Quasars Catalogue). Furthermore, later classifier implementation tests revealed that the exclusion of “Gaia DR3 high-purity QSO” labelled greatly improved the classifier recall. We therefore excluded labels from the “Gaia DR3 high-purity QSO” sample from our final implementation. Further evidence that this sample suffers from contamination by YSOs is discussed in Sect. 5.2.3. A final parent sample with 23 217 and 94 248 sources labelled as extragalactic and stellar was available for the Extragalactic classifier.
Giant classifier sample:
Sources in the Giants sample in Sect. E.2.2 were labelled as such (positive label), and extragalactic sources in Sect. E.2.1 or in the “certainly young” and “other stellar” samples were labelled as “other” (negative label). We deliberately avoid the sample at Orion’s distance as this may contain Giants with unavailable measurements. A final parent sample with 12 988 and 133 966 sources labelled as giant and others was available for the Giant classifier.
The final labelled datasets were split into 30% training, 40% validation and 30% testing sets. We ensured samples contained a representative fraction of the different source types (e.g., resolved and unresolved galaxies) by stratifying them based on origin catalogues (Sect. E.2.1), maintaining similar source proportions from each catalogue. Similarly, each subtype of stellar sources in Sect. E.2.2 was also adequately represented.
E.3 Features
Implementation tests were carried out by training classifiers in six subsets of features, combining available colours from Gaia DR3, PanSTARRS1-DR2, and CatWISE. We included both pure-colour subsets and versions with Gaia parallaxes and uncertainties. A summary of features used in each of our best-performing models is presented in Tab. .
| Model | Classifier | # features | features | sources |
[…]
| Gaia | Giant | 5 | , , , parallax, parallax_error | 8 385 766 |
E.4 Training
Class imbalance was intentionally maintained for both classifiers, as we expect giants and extragalactic sources to be a minority of sources in the OSFC field. However, this choice led to true negative labels dominating precision and accuracy. To reduce its impact on the smaller classes (true positive labels), we prioritised the use of scores that account for the trade-off between precision and recall for model comparison and performance evaluation. Additionally, as a general rule, models with validation sample true positive recovery rate under 75% were discarded. Grid search with 5-fold cross-validation was performed to optimise hyperparameters, using average precision as the reference score. Each classifier was then trained with the optimal parameters and exported.
E.5 Validation
Validation samples were used for the comparison of the performance of each classifier trained, and a series of scores (namely, accuracy, precision, recall, F1-score and F) score) was used to rank their performance. For the extragalactic classifier, we found that sets of attributes covering colours from optical to mid-infrared (GaiaPS1WISE_pure_colour) performed better, qualitatively consistent with results in previously published classifiers. For the giant classifier, the models using Gaia parallax performed better, with the Gaia model (attributes: Gaia colours plus parallax and its uncertainties) being only marginally outperformed by the GaiaWISE model (same Gaia attributes with the addition of CatWISE-based colours). Nevertheless, the requirement for both Gaia and CatWISE data for the latter makes it applicable to a smaller number of sources. We therefore opt to keep the results for the pure-Gaia model (Tab. ).
Validation samples were also used to study the precision-recall curves of the best-performing classifiers. At this stage, we used the F1-score for determining the best classification thresholds while ensuring a good trade-off between precision and recall. Best classification thresholds were found at 0.41 for the extragalactic classifier, and at 0.39 for the giant classifier.
E.6 Classification
We used the classification thresholds from the previous section for classification. Classification reports for the testing set are presented in Tab. 17, and confusion matrices in Fig. 17.
| label | precision | recall | F1-score |
| Extragalactic classifier | |||
| stellar | 1.00 | 1.00 | 1.00 |
| extragalactic | 0.99 | 0.97 | 0.98 |
| Giant Classifier | |||
| Other | 0.99 | 0.99 | 0.99 |
| Giants | 0.94 | 0.93 | 0.93 |
Extragalactic classifier:
We were able to apply the GaiaPS1WISE_pure_colour model to classify 6 520 467 sources, with 139 668 classified as probable extragalactic sources. We note that among stellar sources misclassified as extragalactic (0.19 false negative rate), were spectroscopically confirmed YSOs, although these correspond to only of the YSOs in the testing sample.
Giant Classifier:
We were able to apply the Gaia model to derive the probabilities of being a giant for 8 385 766 sources in the field of the OSFC, with 141 808 likely giants identified. Among sources misclassified as giants in the testing sample, 97% were labelled “other stellar sources“ based on their measurements from spectroscopic surveys. The remaining 3% were real YSOs with spectroscopic constraints, these correspond to of the YSO sources in the testing sample.
The development of these classifiers included 53 154 labelled sources without enough data for classification. Both samples of classified and labelled but not classified sources were matched to the OSFC YSO catalogue and further discussed in Sect. 5.2.